### v1

# Heights and periods of algebraic cycles in families

Abstract.
We consider the Beilinson–Bloch heights and Abel–Jacobian periods of homologically trivial Chow cycles in families. For the Beilinson–Bloch heights, we show that for any $g\geq 3$, there is a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$, the coarse moduli of curves of genus $g$ over $\overline{\mathbb{Q}}$, such that the heights of Ceresa cycles and Gross–Schoen cycles over $U$ satisfy the Northcott property. For the Abel–Jacobi periods, we provide an algebraic criterion for the existence of a Zariski open dense subset of any family such that all cycles not defined over $\overline{\mathbb{Q}}$ are non-torsion and verify that this criterion holds true for Ceresa cycles and Gross–Schoen cycles.
1. Introduction
This paper aims to initiate a project to study Beilinson–Bloch heights and Abel–Jacobi periods for homologically trivial cycles in families. We have achieved two goals: the Northcott property for the heights of the Gross–Schoen cycles and Cerasa cycles and an algebraic criterion for the non-degeneracy of Abel–Jacobi periods for the general family of cycles.
1.1. Gross–Schoen and Ceresa cycles
For a smooth projective irreducible curve $C$ defined over a filed $k$ of $g\geq 3$ and a class $\xi\in\mathrm{Pic}^{1}(C)$ such that $(2g-2)\xi=\omega_{C}$, we have two homologically trivial 1-cycles:
  1. (1)
    the Gross–Schoen cycle ${\mathrm{GS}}(C):=\Delta_{\xi}(C)\in{\mathrm{Ch}}_{1}(C^{3})$ obtained by modified the diagonal cycle in $C^{3}$ using base class $\xi$ ([GS95, Zha10]);
  2. (2)
    the Ceresa cycle by ${\mathrm{Ce}}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in{\mathrm{Ch}}_{1}({\mathrm{Jac}}(C))$ defined by embedding $i_{\xi}\colon C{\longrightarrow}{\mathrm{Jac}}(C)$ via $\xi$.
Up to torsions, the definition of these cycles does not depend on the choice of $\xi$.
Let ${\mathscr{M}}_{g}$ be the moduli space of smooth projective curves of genus $g$. For each $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, denote by ${\mathscr{C}}_{s}$ the curve parametrized by $s$.
Theorem 1.1.
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ defined over ${\mathbb{Q}}$, and positive numbers $\epsilon$ and $c$, such that for any $s\in U(\overline{\mathbb{Q}})$,
$\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
$\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}\geq\epsilon h_{\mathrm{Fal}}(s)-c,$
where $h_{\mathrm{Fal}}$ is the Faltings height on ${\mathscr{M}}_{g}(\overline{\mathbb{Q}})$.
Theorem 1.1 yields the following Northcott property immediately.
Corollary 1.2 (Northcott property).
For each $g\geq 3$, there exists a Zariski open dense subset of the ${\mathscr{M}}_{g}$ defined over $\overline{\mathbb{Q}}$ such that for any $H,D\in{\mathbb{R}}$,
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty,$
#$\#\{s\in U(\overline{\mathbb{Q}}):\quad\deg[{\mathbb{Q}}(s):{\mathbb{Q}}]<D,\quad\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}<H\}<\infty.$
In the course of the proof, we also establish the following geometric result.
Theorem 1.3.
For each $g\geq 3$, there exists a Zariski open dense subset $U$ of ${\mathscr{M}}_{g}$ such that the followings hold true:
  1. (i)
    ${\mathrm{GS}}({\mathscr{C}}_{s})$ and ${\mathrm{Ce}}({\mathscr{C}}_{s})$ are both non-torsion in the Chow groups for all $s\in U({\mathbb{C}})\setminus U(\overline{\mathbb{Q}})$.
  2. (ii)
    there exist at most countably many $s\in U({\mathbb{C}})$ such that ${\mathrm{AJ}}({\mathrm{GS}}({\mathscr{C}}_{s}))$ or ${\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s}))$ is torsion in the intermediate Jacobians.
Hain [Ha24] also proved Theorem 1.3 with $U$ a non-empty analytic open subset of ${\mathscr{M}}_{g}^{\mathrm{an}}$ with a different method. When $g=3$, Hain’s result is completely explicit, while ours is not.
The proofs of Theorem 1.1, Corollary 1.2 and Theorem 1.3 have two parts: the arithmetic part and the geometric part. Both parts are needed to prove Theorem 1.1 and Corollary 1.2, while Theorem 1.3 only relies on the geometric part. We will explain each part in more detail in §$\S$1.3.
Before moving on, let us mention that the geometric part of our paper works for any family of homologically trivial cycles and, more generally, for any admissible normal function. More specifically, for any smooth projective morphism $f\colon X\rightarrow S$ of algebraic varieties with irreducible fibers and any family of homologically trivial cycles $Z$, all defined over $\overline{\mathbb{Q}}$, we show that there exists a Zariski closed subset $S_{{\mathscr{F}}}(1)$ of $S$ such that: (i) $[Z_{s}]$ is non-torsion in ${\mathrm{Ch}}^{*}(X_{s})$ for every transcendental point $s$ of $S\setminus S_{{\mathscr{F}}}(1)$; (ii) there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that ${\mathrm{AJ}}(Z_{s})$ is torsion in the intermediate Jacobian. We then prove a checkable criterion for $S_{{\mathscr{F}}}(1)\not=S$. A simple case will be presented in Corollary 1.9 and Remark 1.10. We do these by studying the normal function associated with $Z$, which is defined in §$\S$§C.5, and by studying its Betti rank.
1.2. Normal functions
Now, we turn to the geometric part of the framework and explain our results.
Let $S$ be a quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ be a VHS of weight $-1$ over $S$, and consider the intermediate Jacobian (write ${\mathscr{V}}$ for the holomorphic vector associated with ${\mathbb{V}}_{{\mathbb{Z}}}$)
$\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})={\mathbb{V}}_{\mathbb{Z}}\backslash{\mathscr{V}}/{\mathscr{F}}^{0}{\mathscr{V}}{\longrightarrow}S.$
Let $\nu\colon S{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function; then $\nu$ defines an admissible variation of mixed Hodge structures ${\mathbb{E}}_{\nu}$ on $S$ which is an extension of ${\mathbb{Z}}(0)_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$.
The fiberwise isomorphism ${\mathbb{V}}_{{\mathbb{R}},s}{\overset{\sim}{{\longrightarrow}}}{\mathbb{V}}_{{\mathbb{C}},s}/{\mathscr{F}}^{0}_{s}={\mathscr{V}}_{s}/{\mathscr{F}}^{0}_{s}$ makes ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ into a local system of real tori
${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{\sim}{\mathbb{V}}_{\mathbb{R}}/{\mathbb{V}}_{\mathbb{Z}}.$
Let ${\mathscr{F}}_{\mathrm{Betti}}$ denote the induced foliation, which we call the Betti foliation. More precisely, for any point $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, there is a local section $\sigma\colon U{\longrightarrow}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ from a neighborhood $\Delta$ of $\pi(x)$ in $S^{\mathrm{an}}$, with $x\in\sigma(U)$, represented by a flat section of ${\mathbb{V}}_{\mathbb{R}}$. The manifolds $\sigma(U)$ gluing together to a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. See §$\S$§C.2 for more details. In particular, all torsion multi-sections are leaves of ${\mathscr{F}}_{\mathrm{Betti}}$.
1.2.1. Betti strata
The Betti foliation defines a strata on $S$ as follows. For each integer $t\geq 0$, set
(1.1) $S_{\mathscr{F}}(t):=\left\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\right\}$
where by abuse of notation $\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}}$ means the intersection with the leaves. This subset is, by definition, real-analytic. We then have the following Betti strata on $S$
(1.2) $\emptyset=S_{{\mathscr{F}}}(\dim S+1)\subseteq S_{{\mathscr{F}}}(\dim S)\subseteq\cdots\subseteq S_{{\mathscr{F}}}(1)\subseteq S_{{\mathscr{F}}}(0)=S.$
A main theorem of the geometric part is that the Betti strata are algebraic:
Theorem 1.4 (Theorem 3.2).
For each $t\geq 0$, $S_{\mathscr{F}}(t)$ is Zariski closed in $S$.
Remark 1.5.
In the Betti strata, $S_{{\mathscr{F}}}(1)$ plays a particularly important role. For example, by definition of the Betti foliation, $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. In particular, there are at most countably many $s\in S({\mathbb{C}})$ outside $S_{{\mathscr{F}}}(1)$ such that $\nu(s)$ is torsion.
Finally, let us point out that $S_{{\mathscr{F}}}(t)$ is closely related to the degeneracy loci defined by the first-named author. Indeed, when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is an abelian scheme, then $S_{{\mathscr{F}}}(t)$ is the $\nu(S)^{\mathrm{deg}}(-t+1)$ in [Gao20a, Defn. 1.6] for each $t\geq 1$.
1.2.2. Betti rank
Now assume that $S$ is smooth. Then the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ for each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. Thus for each $s\in S({\mathbb{C}})$ we have a linear map
(1.3) $\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}.$
The Betti rank of $\nu$ is defined to be:
(1.4) $r(\nu):=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
A trivial upper bound for $r(\nu)$ is $r(\nu)\leq\min\{\dim S,\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for any $s\in S({\mathbb{C}})$. One can also easily improve this trivial upper bound as will be explained below.
The Betti rank is easily seen to be related to the Betti strata in the following way: $\dim S-r(\nu)$ is the maximum of $t\geq 0$ such that $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. Thus Theorem 1.4 implies that $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$. However, this equality is, in general, not applicable to compute $r(\nu)$. But at this stage, we have
(1.5) $S_{{\mathscr{F}}}(1)\not=S\Leftrightarrow r(\nu)=\dim S.$
The second main theorem for the geometric part is the following formula for $r(\nu)$, which is often computable in practice. The VMHS ${\mathbb{E}}_{\nu}$ on $S$ induces a period map $\varphi=\varphi_{\nu}\colon S{\longrightarrow}\Gamma\backslash{\mathscr{D}}$, with ${\mathscr{D}}$ a mixed Mumford–Tate domain with Mumford–Tate group $\mathbf{G}$ (so ${\mathscr{D}}$ is a $\mathbf{G}({\mathbb{R}})^{+}$-orbit for $\mathbf{G}$ the generic Mumford–Tate group of the VMHS ${\mathbb{E}}_{\nu}$); we refer to §$\S$2.12.2 for the construction of $\varphi$ and ${\mathscr{D}}$.
Denote by $V$ the unipotent radical of $\mathbf{G}$. It equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ becomes torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$; see Remark 2.2. Now the trivial upper bound on $r(\nu)$ can be easily improved to be $r(\nu)\leq\min\{\dim\varphi(S),\frac{1}{2}\dim V\}$.
Here is our formula to compute $r(\nu)$.
Theorem 1.6 (Theorem 3.1).
The Betti rank is given by
(1.6) $r(\nu)=\min_{N}\left(\dim\varphi_{/N}(S)+\frac{1}{2}\dim_{\mathbb{Q}}(V\cap N)\right),$
where $N$ runs through the set of normal subgroup of $\mathbf{G}$, and $\varphi_{/N}$ is the induced period map
[…]
with $[p_{N}]$ the quotient by $N$ (see §$\S$§B.3).
Notice that the trivial upper bound on $r(\nu)$ above is recovered by taking $N=\{1\}$ and $N=\mathbf{G}$.
Here are two applications of Theorem 1.6, on two cases where the trivial upper bound on $r(\nu)$ is attained.
Corollary 1.7 (Theorem 5.1).
Assume: (i) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, i.e. the only sub-VHSs are trivial or itself; (ii) $\nu(S)$ is not a torsion section. Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}{\mathbb{V}}_{{\mathbb{Q}},s}\}$ for one (and hence all) $s\in S({\mathbb{C}})$.
Notice that in the situation of Corollary 1.7, we have $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Corollary 1.8 (Theorem 5.2).
Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e. locally constant VHS. Then $r(\nu)=\min\{\dim\varphi(S),\frac{1}{2}\dim_{{\mathbb{Q}}}V\}$.
An immediate corollary on the torsion locus, combining (1.5) and Remark 1.5 and Theorem 1.4, is:
Corollary 1.9.
Under the assumptions of either Corollary 1.7 or Corollary 1.8. Assume that $\varphi$ is generically finite and that $\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V$. Then there exists a Zariski open dense subset $U$ of $S$ such that $\nu(s)$ is torsion for at most countably many $s\in U({\mathbb{C}})$. Indeed, one can take $U=S\setminus S_{{\mathscr{F}}}(1)$.
Remark 1.10.
If $\varphi$ is not necessarily generically finite, then the conclusion becomes: there exists a Zariski open dense subset $U$ of $S$ such that $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is contained in at most countably many fibers of $\varphi$. Indeed, $\varphi(S)$ is an algebraic variety by [BBT23]. Then we can conclude by applying Corollary 1.9 to $\varphi(S)$ and the induced normal function $\nu^{\prime}$ (see §$\S$3.1 for the construction of $\nu^{\prime}$).
Applying either Corollary 1.7 or Corollary 1.8 to the Gross–Schoen and the Ceresa normal functions, we obtain the following corollary will be used to prove Theorem 1.1, Corollary 1.2 and Theorem 1.3.
Corollary 1.11 (Corollary 5.3).
If $g\geq 3$, then the Betti rank of the Gross-Schoen (resp. of the Ceresa) normal function is $33g-$ 3.
Hain [Ha24] gave a different proof of Corollary 1.11. When $g=3$, he even gave an explicit description on the locus where $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<3g-3$.
1.3. Plan and Ingredients of proofs
Our plan to prove results like Theorem 1.1, Corollary 1.2 and Theorem 1.3 has two parts: the arithmetic part and the geometric part. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers and let $Z$ be a family of homologically trivial cycles, all defined over $\overline{\mathbb{Q}}$
The arithmetic part requires the construction of a suitable adelic line bundle $\overline{{\mathscr{L}}}$ over the base, a theory initiated by the second-named author and developed in further joint work with Yuan [YZ21], such that the height function $h_{\overline{{\mathscr{L}}}}$ is the Beilinsin–Bloch height. Then, the desired height inequalities will follow from suitable bigness properties of $\overline{{\mathscr{L}}}$. In the context of Theorem 1.1 and Corollary 1.2, by the second-named author’s [Zha10, Thm. 2.5.5] the desired $\overline{{\mathscr{L}}}$ can be obtained from a suitable Deligne pairing. The required bigness property is the bigness of the generic fiber $\tilde{{\mathscr{L}}}$ of $\overline{{\mathscr{L}}}$. We furthermore show in Proposition 6.3 that the bigness of $\tilde{{\mathscr{L}}}$ will follow from the non-vanishing of $c_{1}(\overline{{\mathscr{L}}})^{\wedge\dim S}$, with $c_{1}(\overline{{\mathscr{L}}})$ the curvature form. The arithmetic part of this paper is confined to the Gross-Shoen and the Ceresa cycles.
The geometric part studies the (admissible) normal function $\nu$ associated with $Z$. The main results are the Zariski closedness of the Betti strata and the formula for the Betti rank $r(\nu)$. This part has been explained in §$\S$1.2. The proofs of Theorem 1.4 and Theorem 1.6 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the generic rank of the Betti map for abelian schemes. A core of our proof is Ax–Schanuel for VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24], which will be used multiple times.
Hain’s works on the Hodge-theoretic computation of the Archimedean local height pairing and the Betti form are the key ingredients to bridging the geometric and arithmetic parts. Indeed, Hain in [Hai90] proved that the archimedean local height pairing could be computed using the metricized biextension line bundle on $S$, and in [HR04] computed (joint with Reed), the curvature form of the metrized biextension line bundle. In our situation, we work with the height pairing of $Z_{s}$ with itself for any $s\in S(\overline{\mathbb{Q}})$, and the metrized biextension line bundle in question is the pullback of the metrized tautological line bundle under $\nu$. Hain in [Hai13] proved that its curvature form $\beta_{\nu}$, called the Betti form, is a semi-positive $(1,1)$-form. We show in Corollary D.7 that $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $r(\nu)=\dim S$. Finally, back to the situation of Theorem 1.1, Corollary 1.2, and Theorem 1.3. Hain’s formula [Hai90, Prop. 3.3.12] and the second-named author’s [Zha10, Thm. 2.5.5] together imply that the curvature $c_{1}(\overline{{\mathscr{L}}})$ is precisely the Betti form. This finishes the bridge between the geometric part and the arithmetic part. Again, the bridge between the geometric part and the arithmetic part works in much larger generality, provided that the arithmetic part is solved.
Acknowledgements
We would like to thank Richard Hain for numerous inspiring discussions and for sharing his notes and Xinyi Yuan for his help during the preparation of this project. Part of this work was done when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. Finally, we would like to thank Richard Hain for mentioning our work and method at the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM in May, 2024.
Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714). Shouwu Zhang is supported by an NSF award, DMS-2101787.
2. Period map associated with normal functions
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function.
The first goal of this section is to define the period map $\varphi_{\nu}$ associated with $\nu$ (when $\nu$ is clearly in context, we simply denote it by $\varphi$). This map fits into the diagram in the category of complex varieties
(2.1) ${{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\Gamma\backslash{\mathscr{D}}}$
where
  1. (i)
[…]
  2. (ii)
    $\Gamma$ is a suitable arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$.
We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$, and its geometric meaning will be given in Remark 2.2.
[…]
2.1. Universal period map to the classifying space
Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$ with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
Let $u_{S}\colon\tilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map. Then the pullback $u_{S}^{*}{\mathbb{E}}_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\cong\tilde{S}\times E_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\cong\tilde{S}\times M_{-1,{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp. with $M_{-1,{\mathbb{Z}}}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$). Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $M_{-1,{\mathbb{Z}}}$. The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\tilde{S}$ is simply connected.
Each $\tilde{s}\in\tilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\tilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\tilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map
$\tilde{\varphi}=\tilde{\varphi}_{\nu}\colon\tilde{S}\longrightarrow{\mathscr{M}}$
[…]
(2.2) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\tilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varphi}}$$\scriptstyle{\tilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
2.2. Construction of the Mumford–Tate domain
In practice, we need to refine this period map and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$, which contains $\tilde{\varphi}(\tilde{S})$. It is constructed as follows.
[…]
Fix $\tilde{s}\in\tilde{S}$. Define
(2.3) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\tilde{\varphi}(\tilde{s})\subseteq{\mathscr{M}}.$
[…]
Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\tilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$ fitting into the diagram (2.1).
2.3. Setup for o-minimality
Let $\mathfrak{F}\subseteq{\mathscr{D}}$ be a fundamental set for the quotient $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$, i.e. $u|_{\mathfrak{F}}$ is surjective and $(u|_{\mathfrak{F}})^{-1}(\bar{x})$ is finite for each $\bar{x}\in\Gamma\backslash{\mathscr{D}}$. If $\mathfrak{F}$ is a semi-algebraic subset of ${\mathscr{D}}$, then we have a semi-algebraic structure on $\Gamma\backslash{\mathscr{D}}$ induced by $u|_{\mathfrak{F}}$.
By the main result of [BBKT24], there exists a semi-algebraic fundamental set $\mathfrak{F}\subseteq{\mathscr{D}}$ for the quotient $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ with the following properties: $\varphi$ is ${\mathbb{R}}_{\mathrm{an},\exp}$-definable for the semi-algebraic structure on $\Gamma\backslash{\mathscr{D}}$ defined by $\mathfrak{F}$.
2.4. Relating the Betti foliation and the period map
Recall the projection to the pure part $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}\subseteq{\mathscr{M}}_{0}$ from (2.2) and the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i). The diagram (2.1) can be complete into:
[…]
and $\varphi_{0}=p\circ\varphi$ is the period map for the VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow S$.
For each $\tilde{s}\in\tilde{S}$, write $\tilde{s}_{V}$ for the image $\tilde{\varphi}(\tilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$. This gives a map $\tilde{\varphi}_{V}\colon\tilde{S}\rightarrow V({\mathbb{R}})$ sending $\tilde{s}\mapsto\tilde{s}_{V}$. Notice that
(2.5) $\tilde{\varphi}=(\tilde{\varphi}_{V},\tilde{\varphi}_{0})\colon\tilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
Recall $S_{{\mathscr{F}}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define $S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$ for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.3).
Lemma 2.1.
For each $t\geq 0$, we have
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)$ =$=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\{\tilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
(2.6) $=\bigcup\nolimits_{r\geq 0}\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\tilde{\varphi}(\tilde{s}))=r,~{}\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})\text{ for }$
[…]
and
(2.7) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\tilde{s}\in\tilde{S}:\mathrm{rank}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}\leq t\}.$
Notice that in the union in (2.1), the second condition $\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})$ always holds true for $r>t$.
Proof.
Consider $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lies in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ for $s=\pi(x)$, which is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$ by Carlson [Car85]. Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$. The map $\nu\circ u_{S}\colon\tilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\tilde{\nu}\colon\tilde{S}\rightarrow\tilde{{\mathscr{J}}}$, and $\tilde{\varphi}=\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}$.
[…]
The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\tilde{\varphi}_{{\mathscr{J}}}$ becomes
(2.8) $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}\xrightarrow{(1,\tilde{\varphi}_{0})}M_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}.$
By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in M_{-1}({\mathbb{R}})$.
Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S_{{\mathscr{F}}}(t))\right)=\{\tilde{s}^{\prime}\in\tilde{\nu}(\tilde{S}):\dim_{\tilde{s}^{\prime}}\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\tilde{s}^{\prime}_{V}$ is the image of $\tilde{s}\in\tilde{{\mathscr{J}}}\xrightarrow{\tilde{\varphi}_{{\mathscr{J}}}}M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow M_{-1}({\mathbb{R}})$ with the last map being the natural projection. Applying $\tilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\tilde{\nu}=\nu\circ u_{S}$, we have
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}(\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu})^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
Hence the first equality in (2.1) holds true because $\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}=\tilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$. Similarly, we get (2.7).
The second equality in (2.1) clearly holds. Hence, we are done. ∎
Remark 2.2.
The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion. The relative dimension $\dim{\mathscr{J}}_{\nu}-\dim S$ equals $\frac{1}{2}\dim V$.
Similarly, $V$ equals ${\mathbb{V}}^{\prime}_{{\mathbb{Q}},s}$ for any $s\in S({\mathbb{C}})$, where ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}$ is the largest sub-VHS of ${\mathbb{V}}_{{\mathbb{Z}}}$ such that $\nu$ is torsion under the projection ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}/{\mathbb{V}}^{\prime}_{{\mathbb{Z}}})$.
Moreover for the natural projection $\varphi_{{\mathscr{J}}_{\nu}}\colon{\mathscr{J}}_{\nu}\rightarrow\Gamma\backslash{\mathscr{D}}$, we have $\varphi(S)=\varphi_{{\mathscr{J}}_{\nu}}(\nu(S))$.
Since $\tilde{\varphi}_{V}$ factors through $\tilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.7) immediately yields the following trivial upper bound
(2.9) $\nu_{\mathrm{Betti},s}(T_{s}S)\leq\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}\quad\text{ for all }s\in S({\mathbb{C}}).$
3. The Betti rank and Zariski closedness of the rank-strata
The goal of this section is to prove one of the main theorems of this paper. We will prove a formula to compute the Betti rank $r(\nu)$, and in the process, we show that the Betti foliation on the intermediate Jacobian defines Zariski closed strata.
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function.
[…]
$r(\nu)=\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S).$
Retain the notation from (2.1). In particular, we have the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$ for the (mixed) Mumford–Tate domain ${\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group. We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$.
The main result of this section is the following formula for $r(\nu)$. The advantage is that it is often computable in practice.
Theorem 3.1.
The Betti rank is given by the following formula:
(3.1) $r(\nu)=\min_{N}\left\{\dim\varphi_{/N}(S)+\frac{1}{2}\dim_{{\mathbb{Q}}}(V\cap N)\right\},$
where $N$ runs through the set of normal subgroups of $\mathbf{G}$, and $\varphi_{/N}$ is the induced map
[…]
Taking $N=\{1\}$ and $N=\mathbf{G}$, we recover the trivial upper bound (2.9) of $r(\nu)$.
We also show that the Betti foliation defines Zariski’s closed strata on $S$.
Theorem 3.2.
For each $t\geq 0$, the set $S_{{\mathscr{F}}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$. In particular, $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$.
The proofs of Theorem 3.1 and Theorem 3.2 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the case when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable. A key ingredient for our proof is the mixed Ax-Schanuel theorem, which is used multiple times. In [Gao20a], the version for universal abelian varieties [Gao20b, Thm. 1.1] was used. In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, for two reasons. First, it eases the notation for the proof. Second and more importantly, [BBT23] is necessary to prove the full version of Theorem 1.6. Still, we point out that [BBT23] is not needed in the case which we care the most about in this paper, i.e. $r(\nu)=\dim S$ if and only if the RHS of (1.6) equals $\dim S$ because we can easily reduce the proof to the case where every fiber of $\varphi$ has dimension 00.
Another crucial input for the proof is the o-minimal structure associated with the period map [BBKT24], which we will recall in §$\S$2.3. This allows us to apply (o-minimal) definable Chow.
3.1. Replacing $S$ by $\varphi(S)$
We shall replace $S$ by $\varphi(S)$ in the proof using [BBT23]. This largely eases the notation.
[…]
${S^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Gamma\backslash{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[p]}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{0}}$$\scriptstyle{\varphi}$${\Gamma_{0}\backslash{\mathscr{D}}_{0}}$
with the dotted arrow being the restriction $[p]|_{S^{\prime}}$.
Recall from Remark 2.2 that ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$ is an intermediate Jacobian over $S$. Set ${\mathscr{J}}^{\prime}:=S^{\prime}\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})$; it is an intermediate Jacobian over $S^{\prime}$. Then the inclusion $S^{\prime}\subseteq\Gamma\backslash{\mathscr{D}}$ yields a section $\nu^{\prime}$ of ${\mathscr{J}}^{\prime}\rightarrow S^{\prime}$, and thus we can define $S^{\prime}_{{\mathscr{F}}}(t)$ for each $t\geq 0$ with respect to $\nu^{\prime}$. We have the following commutative diagram, with $\varphi_{{\mathscr{J}}}$ induced by $\varphi$, such that $\varphi_{{\mathscr{J}}}\circ\nu=\nu^{\prime}\circ\varphi$:
[…]
For each $r\geq 0$, denote by
(3.2) $S_{\geq r}:=\{s\in S({\mathbb{C}}):\dim_{s}\varphi^{-1}(\varphi(s))\geq r\}.$
It is a closed algebraic subset of $S$ by upper semi-continuity.
By (the proof of) Lemma 2.1, more precisely the second equality of (2.1), we have
(3.3) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
This equality allows us to replace $S$ by $\varphi(S)$ to study the Betti rank and the Betti strata.
3.2. Bi-algebraic system and Ax–Schanuel
From now on, in the whole section, we replace $S$ by $\varphi(S)$ and view $S$ as an algebraic subvariety of the complex analytic space $\Gamma\backslash{\mathscr{D}}$, unless otherwise stated.
[…]
Proposition 3.3.
Let ${\mathscr{D}}_{N}$ be a weak Mumford–Tate domain. Then $u({\mathscr{D}}_{N})\cap S$ is a closed algebraic subset of $S$.
Definition 3.4.
  1. (i)
    Let $\tilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset. The weakly special closure of $\tilde{Y}$, denoted by $\tilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\tilde{Y}$.
  2. (ii)
    Let $Y\subseteq S$ be an irreducible subvariety. The weakly special closure of $Y$, denoted by $Y^{\mathrm{ws}}$, is $u(\widetilde{Y}^{\mathrm{ws}})$ for one (hence any) complex analytic irreducible component $\widetilde{Y}$ of $u^{-1}(Y)$.
The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
Theorem 3.5 (weak Ax–Schanuel for VMHS).
Let $\tilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
(3.4) $\dim\tilde{Z}^{\mathrm{Zar}}+\dim u(\tilde{Z})^{\mathrm{Zar}}\geq\dim\tilde{Z}^{\mathrm{ws}}+\dim\tilde{Z},$
where $\tilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\tilde{Z}$.
Proof.
Let $Y:=u(\tilde{Z})^{\mathrm{Zar}}$. Let $\mathscr{Z}:=\{(z,y)\in\tilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$. The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\tilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\tilde{Z}$. Then (3.4) is a direct consequence of the mixed Ax–Schanuel theorem [GK24, Thm. 1.1] applied to $\mathscr{Z}$. ∎
We close this introductory subsection with the following definition. In practice, we often need to work with algebraic subvarieties $Y\subseteq S$, which are not weak Mumford–Tate domains, and the following number measures how far it is from being one.
(3.5) $\delta_{\mathrm{ws}}(Y):=\dim Y^{\mathrm{ws}}-\dim Y.$
If we do not replace $S$ by $\varphi(S)$, then each $Y$ on the right-hand side should be replaced by $\varphi(Y)$.
Definition 3.6.
An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds true: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
3.3. Applications of Ax–Schanuel
Retain the notation in (2.4). We start with the following application of mixed Ax–Schanuel.
Proposition 3.7.
For each $t\geq 0$, $S_{{\mathscr{F}}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
(3.6) $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t.$
Proof.
It suffices to prove two things:
  1. (i)
    $S_{{\mathscr{F}}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
  2. (ii)
    If $Y\subseteq S$ is an irreducible subvariety satisfying (3.6) and is maximal for this property with respect to inclusions, then $Y$ is weakly optimal.
Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S_{{\mathscr{F}}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
$Y:=\overline{u(\{a\}\times\tilde{C})}^{\mathrm{Zar}},$
for some complex analytic irreducible $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}=t$ and some $a\in V({\mathbb{R}})$.
Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\tilde{C}$). Then we get
$\dim\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}+t.$
By Lemma B.4, $\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\tilde{C}}^{\mathrm{Zar}}$. Hence
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\overline{\tilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
The last inequality holds because $\overline{\tilde{C}}^{\mathrm{Zar}}\subseteq\tilde{C}^{\mathrm{ws}}$. So
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
Now, to prove (i), it suffices to prove $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
Let us prove $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$. By definition of $Y$, we have $u(\{a\}\times\tilde{C})\subseteq Y$. Hence $u((\{a\}\times\tilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$. On the other hand, $u((\{a\}\times\tilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. Now we have established $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
Similarly, we have $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
For (ii), let $Y\subseteq Y^{\prime}\subseteq X$. Assume $\delta_{\mathrm{ws}}(Y)\geq\delta_{\mathrm{ws}}(Y^{\prime})$, i.e.
$\dim Y^{\mathrm{ws}}-\dim Y\geq\dim Y^{\prime,\mathrm{ws}}-\dim Y^{\prime}.$
The assumption on $Y$ implies $\dim Y^{\mathrm{ws}}-\dim Y\leq\dim[p](Y^{\mathrm{ws}})-t$. Combined with the inequality above, we obtain $\dim Y^{\prime,\mathrm{ws}}-\dim Y^{\prime}\leq\dim[p](Y^{\prime,\mathrm{ws}})-t$ because $Y\subseteq Y^{\prime}$. Therefore $Y=Y^{\prime}$ by maximality of $Y$. Hence, (ii) is established. ∎
Next, we state a finiteness proposition for weakly optimal subvarieties in $S$ la Ullmo, which we give a proof using twice mixed Ax–Schanuel in the next section. When ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable, i.e. in the mixed Shimura case, this is [Gao20b, Thm. 1.4].
Proposition 3.8.
There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds true. For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}$ is the image of an $N_{j}({\mathbb{R}})^{+}$-orbit contained in ${\mathscr{D}}^{\prime}_{j}$ under $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ for some $j\in\{1,\ldots,k\}$.
The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds true without replacing $S$ by $\varphi(S)$.
Denote by $\Gamma_{j}=\Gamma\cap\mathrm{MT}({\mathscr{D}}^{\prime}_{j})({\mathbb{Q}})$ and $\Gamma_{j,/N_{j}}=\Gamma_{j}/(\Gamma_{j}\cap N_{j}({\mathbb{Q}}))$. Then equivalently, each such $Y^{\mathrm{ws}}$ is a fiber of the quotient $[p_{N_{j}}]\colon u({\mathscr{D}}^{\prime}_{j})=\Gamma_{j}\backslash{\mathscr{D}}^{\prime}_{j}\rightarrow\Gamma_{j,/N_{j}}\backslash({\mathscr{D}}^{\prime}_{j}/N_{j})$.
[…]
3.4. Proof of Theorem 3.1
For each $j\in\{1,\ldots,k\}$, Proposition 3.3 says that $u({\mathscr{D}}^{\prime}_{j})\cap S$ is a closed algebraic subset of $S$. The restriction
$[p_{N_{j}}]|_{S}\colon u({\mathscr{D}}^{\prime}_{j})\cap S\rightarrow\Gamma_{j,/N_{j}}\backslash({\mathscr{D}}^{\prime}_{j}/N_{j})$
[…]
For each $t\geq 0$, the subset
(3.7) $E_{j}(t):=\left\{s\in u({\mathscr{D}}^{\prime}_{j})\cap S:\dim_{s}[p_{N_{j}}]|_{S}^{-1}([p_{N_{j}}](s))\geq\frac{1}{2}\dim(V\cap N_{j})+t\right\}$
is both definable and complex analytic in $u({\mathscr{D}}^{\prime}_{j})\cap S$. Hence, $E_{j}(t)$ is algebraic by definable Chow. Moreover, it is closed in $u({\mathscr{D}}^{\prime}_{j})\cap S$ by the upper semi-continuity of fiber dimensions. So $E_{j}(t)$ is a closed algebraic subset of $S$.
Proposition 3.9.
For each $t\geq 0$, we have
$S_{{\mathscr{F}}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
Proof.
Let $t\geq 0$. By Proposition 3.7, $S_{{\mathscr{F}}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$. Then by Proposition 3.8, $Y^{\mathrm{ws}}$ is a fiber $[p_{N_{j}}]$ for some $j\in\{1,\ldots,k\}$, and hence $\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})=\frac{1}{2}\dim(V\cap N_{j})$. So $\dim Y\geq\frac{1}{2}\dim(V\cap N_{j})+t$. So $Y\subseteq E_{j}(t)$ because $[p_{N_{j}}](Y)$ is a point. ∎
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
[…]
Let us prove $\leq$”. By (2.7) we have $r(\nu)=\max_{\tilde{s}\in\tilde{S}}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}$. Hence $r(\nu)\leq\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)$ for any normal subgroup $N\lhd\mathbf{G}$.
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
$S_{{\mathscr{F}}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
Each $S_{\geq r}$ is Zariski closed in $S$, and each $E_{j}(t-r)$ is Zariski closed in $S^{\prime}=\varphi(S)$. Hence, each member in the union on the right-hand side is Zariski closed in $S$. Taking the Zariski closure of both sides, we then have $S$ equal to a member on the right-hand side.
If $S=S_{\geq t}$, then $\geq$” holds true already for $N=\{1\}$.
Assume $S=S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right)$ for some $10\leq r\leq t-$ 1 and some $j$. Then $S=S_{\geq r}=\varphi^{-1}\left(E_{j}(t-r)\right)$, So each fiber of $\varphi$ has dimension $\geq r$, and $S^{\prime}=E_{j}(t-r)$. Moreover, $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})=\mathbf{G}$ because $S^{\prime}=\varphi(S)$ is Hodge generic in $\Gamma\backslash{\mathscr{D}}$. Set $N=N_{j}$. Each fiber of the map
[…]
has ${\mathbb{C}}$-dimension $\geq r+\left(\frac{1}{2}\dim(V\cap N)+(t-r)\right)=\frac{1}{2}\dim(V\cap N)+t$ by definition of $E_{j}(t-r)$. So
$r(\nu)=\dim S-t\geq\left(\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)+t\right)-t=\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N).$
So, $\geq$” is established. ∎
3.5. Zariski closedness of the degeneracy loci
We start with the following lemma, which is the converse of Proposition 3.9.
Lemma 3.10.
For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S_{{\mathscr{F}}}(t)$.
Proof.
Fix $j$. Denote by $\mathbf{H}_{j}=\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, $V_{j}:=V\cap\mathbf{H}_{j}$, and $\mathbf{H}_{j,0}:=\mathbf{H}_{j}/V_{j}$. Under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}_{j}=(V_{j}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime}_{j})$ by Lemma B.5 (applied to ${\mathscr{D}}^{\prime}_{j}\subseteq{\mathscr{D}}$).
Because $N_{j}\lhd\mathbf{H}_{j}$, we have: (i) $V\cap N_{j}=V_{j}\cap N_{j}$ is a $\mathbf{H}_{j,0}$-module; (ii) the action of $p(N_{j})\lhd\mathbf{H}_{j,0}$ on $V_{j}/(V_{j}\cap N_{j})$ is trivial. Let $x\in{\mathscr{D}}^{\prime}_{j}$. Under ${\mathscr{D}}^{\prime}_{j}=(V_{j}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime}_{j})$, write $x=(v,x_{0})$. Then $N_{j}({\mathbb{R}})^{+}x$ becomes $((V\cap N_{j})({\mathbb{R}})+v)\times p(N_{j})({\mathbb{R}})^{+}x_{0}$. Notice that this $v\in V({\mathbb{R}})$ is fixed.
For each $s\in E_{j}(t)$, by definition there exist an irreducible $\tilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\tilde{Y})$, $\dim\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\tilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$. The last condition implies that $\tilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$. Hence by the discussion above, $\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$. Now that $\dim_{{\mathbb{C}}}\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds true: For each $(a,x_{0})\in\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}\geq t$ such that $\{a\}\times\tilde{C}\subseteq\tilde{Y}$. Hence $s\in S_{{\mathscr{F}}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)). Now, the conclusion of the lemma holds as $s$ runs over $E_{j}(t)$. ∎
Now we are ready to prove Theorem 3.2.
Proof of Theorem 3.2.
[…]
By (3.3), Proposition 3.9 and Lemma 3.10 (both applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
(3.8) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
So $S_{{\mathscr{F}}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is. The “In particular” part is easy to check once we have established the Zariski closedness of $S_{{\mathscr{F}}}(t)$. ∎
4. Proof of the finiteness result à la Ullmo
The goal of this section is to prove Proposition 3.8, the finiteness result regarding weakly optimal subvarieties for admissible VMHS of weight $-1$ and 00.
Before moving on to the proof, let us take a step back and look at this proposition from a historical point of view. In studying the André–Oort conjecture, Ullmo [Ull14, Thm. 4.1] proved this finiteness result for maximal weakly special subvarieties (a particular kind of weakly optimal subvarieties) in the case of pure Shimura varieties as an application of the pure Ax–Lindemann theorem (a special case of Ax–Schanuel). The finiteness is ultimately obtained by the following fact: any countable set which is definable in an o-minimal structure is finite.
[…]
Recently, Baldi–Urbanik [BU24, Thm. 7.1] proved Proposition 3.8 for general admissible VMHS in a different way (but also uses Ax–Schanuel as a core). Weakly optimal subvarieties are called monodromically atypical maximal, and Proposition 3.8 is called Geometric Zilber–Pink as in [BKU24]. Their proof does not use o-minimality, and gives some effective results.
Retain the notation in (2.4). In this section, we also replace $S$ by $\varphi(S)$ to ease notation. The proof also works without doing so, except that when we apply Ax–Schanuel for Theorem 4.2, we need to combine it with [BBT23].
4.1. Zariski optimal subsets and an application of mixed Ax–Schanuel
Definition 4.1.
A complex analytic irreducible subset $\tilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds true: $\tilde{Y}\subsetneq\tilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$, with $\tilde{Y}^{\prime}$ complex analytic irreducible.
The following theorem is, in fact, an equivalent statement to the weak mixed Ax–Schanuel theorem (Theorem 3.5). In this paper, we only need one deduction.
Theorem 4.2.
Assume $\tilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ and $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
Proof.
Let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal. Then $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$. Hence it suffices to prove $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\tilde{Y}$. Then $u(\tilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\tilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
Assume $\tilde{Y}\not=\tilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$ by the Zariski optimality of $\tilde{Y}$. So
$\dim\tilde{Y}^{\mathrm{Zar}}-\dim\tilde{Y}<\dim\tilde{Y}^{\prime\mathrm{Zar}}-\dim\tilde{Y}^{\prime}\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Now that $u(\tilde{Y})\subseteq u(\tilde{Y}^{\prime})$, we have $u(\tilde{Y})^{\mathrm{Zar}}\subseteq u(\tilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
$\dim u(\tilde{Y})^{\mathrm{Zar}}\leq\dim u(\tilde{Y}^{\prime})=\dim\tilde{Y}^{\prime}.$
These two inequalities together yield.
$\dim\tilde{Y}^{\mathrm{Zar}}+\dim u(\tilde{Y})^{\mathrm{Zar}}<\dim\tilde{Y}^{\mathrm{ws}}+\dim\tilde{Y}.$
This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\tilde{Y}$.
Hence we must have $\tilde{Y}=\tilde{Y}^{\prime}$. In particular, $u(\tilde{Y})$ is algebraic. So $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$. ∎
For $Y\subseteq S$ irreducible, let $\tilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
Corollary 4.3.
If $Y\subseteq X$ is weakly optimal, then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
Proof.
Let $\tilde{Y}^{\prime}\supseteq\tilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\tilde{Y})\geq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}^{\prime}$ is Zariski optimal.
Set $Y^{\prime\prime}:=u(\tilde{Y}^{\prime})^{\mathrm{Zar}}$. Then $Y^{\prime\prime}\subseteq u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic. Thus
$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\tilde{Y}^{\prime,\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Since $\tilde{Y}^{\prime}$ is Zariski optimal, we have $\tilde{Y}^{\prime,\mathrm{Zar}}=\tilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\tilde{Y}^{\prime,\mathrm{Zar}}-\dim\tilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\tilde{Y})\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}=\delta_{\mathrm{ws}}(Y).$
So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\tilde{Y}=\tilde{Y}^{\prime}$ is Zariski optimal. ∎
4.2. A parametrization of Zariski optimal subsets
The goal of this subsection is to construct a space ${\mathscr{N}}^{0}$ which parametrizes Zariski optimal subsets of $u^{-1}(S)$.
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by $(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0})$.
[…]
Consider the set ${\mathscr{N}}$ of pairs $(x,H)$ consisting of $x\in(u|_{\mathfrak{F}})^{-1}(S)$ and $H$ a connected subgroup of $\mathbf{G}_{{\mathbb{R}}}$ with the following properties: $H_{0}:=H/(V_{{\mathbb{R}}}\cap H)$ is a semi-simple subgroup of $\mathbf{G}_{0,{\mathbb{R}}}$, and $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$.
For each $(x,H)\in{\mathscr{N}}$, denote by $V_{H}:=V_{{\mathbb{R}}}\cap H$. Then $V_{H}$ is the unipotent radical of $H$. It can be easily shown that each weak Mumford–Tate domain is $N({\mathbb{R}})x$ for some $(x,N_{{\mathbb{R}}})\in{\mathscr{N}}$ (with $N_{{\mathbb{R}}}$ defined over ${\mathbb{Q}}$).
Lemma 4.4.
Let $(x,H)\in{\mathscr{N}}$. Then $H({\mathbb{R}})x$ is:
  1. (i)
    an algebraic subset of ${\mathscr{D}}$;
  2. (ii)
    $(v^{\prime}+V_{H}({\mathbb{R}}))\times H_{0}({\mathbb{R}})x_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}$ for some $v^{\prime}\in V({\mathbb{R}})$ and $x_{0}=p(x)$.
Proof.
$H({\mathbb{R}})x$ is by definition semi-algebraic, and is complex analytic since $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$. So (i) holds.
Any two Levi decompositions differ from the conjugation by an element of the unipotent radical. Hence there exists a $v\in V({\mathbb{R}})$ such that
$H=(v,1)(V_{H}\rtimes H_{0})(-v,1).$
Thus the condition $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(H)$ implies $\mathrm{Int}(-v)(h_{x}({\mathbb{S}}))\subseteq\mathrm{N}_{\mathbf{G}_{{\mathbb{R}}}}(V_{H}\rtimes H_{0})=V^{\prime}\rtimes G_{0}^{\prime}$ for some $V^{\prime}<V_{{\mathbb{R}}}$ and $G_{0}^{\prime}<\mathbf{G}_{0,{\mathbb{R}}}$. Then $H_{0}$ acts trivially on $V^{\prime}/V_{H}$ by the normality condition.
Write $x=(x_{V},x_{0})\in V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}$, then $\mathrm{Int}(-v)(h_{x}({\mathbb{S}}))=h_{(x_{V}-v,x_{0})}({\mathbb{S}})$ by Proposition B.3.(i). So by the last paragraph $x_{V}-v\in V_{H}({\mathbb{R}})+v^{\prime\prime}$ for some $v^{\prime\prime}\in V({\mathbb{R}})$ with $H_{0}({\mathbb{R}})\cdot v^{\prime\prime}=v^{\prime\prime}$. Thus $(v,1)(V_{H}\rtimes H_{0})({\mathbb{R}})(-v,1)\cdot x=(V_{H}({\mathbb{R}})+v^{\prime\prime}+v)\times H_{0}({\mathbb{R}})x_{0}$. So (ii) holds. ∎
Define the following two functions on ${\mathscr{N}}$:
(4.1) $d\colon{\mathscr{N}}\rightarrow{\mathbb{R}},$ $\quad(x,H)\mapsto\dim_{x}\left(u^{-1}(S)\bigcap H({\mathbb{R}})x\right),$
$\delta\colon{\mathscr{N}}\rightarrow{\mathbb{R}},$ $\quad(x,H)\mapsto\dim_{x}H({\mathbb{R}})x-\dim_{x}\left(u^{-1}(S)\bigcap H({\mathbb{R}})x\right).$
Finally, we are ready to define
[…]
$H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime}),$
$H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x\Rightarrow\delta(x,H)<\delta(x,H^{\prime})\big{\}}.$
The proof of the following proposition uses Theorem 4.2 (twice) and hence mixed Ax–Schanuel.
Proposition 4.5.
The following two sets are equal:
    the set of orbits $\{H({\mathbb{R}})x:(x,H)\in{\mathscr{N}}^{0}\}$;
    $\{\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}:\tilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\tilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
Moreover, $H({\mathbb{R}})x$ is a weak Mumford–Tate domain for each $(x,H)\in{\mathscr{N}}^{0}$.
Proof.
Take $(x,H)\in{\mathscr{N}}^{0}$. We wish to prove that $H({\mathbb{R}})x$ equals $\tilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\tilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\tilde{Y}^{\prime}=d(x,H)$.
Take $\tilde{Y}\supseteq\tilde{Y}^{\prime}$ with $\tilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\tilde{Y})\leq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\tilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$. Thus $\delta_{\mathrm{Zar}}(\tilde{Y})=\delta(x,H^{\prime})$.
Thus $\tilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$. Hence $d(x,H)=d(x,H^{\prime\prime})$, so by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime\prime}({\mathbb{R}})x$. So $H({\mathbb{R}})x\subseteq H^{\prime}({\mathbb{R}})x$. But $\delta(x,H)=\delta(x,H^{\prime})$. So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$. Now we have proved that the first set is contained in the second set and the “Moreover” part of the proposition.
Conversely let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. By Theorem 4.2, $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$. We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\tilde{Y}=d(x,H)$.
We wish to prove $(x,H)\in{\mathscr{N}}^{0}$. Let us check the two properties which define ${\mathscr{N}}^{0}$ by contradiction.
Assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x$ and $d(x,H)=d(x,H^{\prime})$. Then $\dim\tilde{Y}=d(x,H^{\prime})$ and hence $\tilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation. But then $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i). So $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$, which is impossible. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime})$.
Next, assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$ and $\delta(x,H)\geq\delta(x,H^{\prime})$. Take a complex analytic irreducible component $\tilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\tilde{Y}^{\prime}=d(x,H^{\prime})$. Then $\tilde{Y}\subseteq\tilde{Y}^{\prime}$ by analytic continuation. By Lemma 4.4.(i), we have $\tilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$. So $\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\tilde{Y})$. Hence $\tilde{Y}^{\prime}=\tilde{Y}$ because $\tilde{Y}$ is Zariski optimal. So $\delta(x,H^{\prime})=\delta(x,H)$ and $d(x,H^{\prime})=d(x,H)$, and so
$\dim H({\mathbb{R}})x=d(x,H)+\delta(x,H)=d(x,H^{\prime})+\delta(x,H^{\prime})=\dim H^{\prime}({\mathbb{R}})x.$
This contradicts $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x\Rightarrow\delta(x,H)<\delta(x,H^{\prime})$.
This finishes the proof of $(x,H)\in{\mathscr{N}}^{0}$. So, the second set is contained in the first set. Now we are done. ∎
4.3. A finiteness result for ${\mathscr{N}}^{0}$
Let ${\mathscr{N}}^{0}$ be as defined in (4.2). In this subsection, we prove that the following finiteness result. The proof relies on o-minimality.
Proposition 4.6.
There are only finitely many subgroups $H$ of $\mathbf{G}_{{\mathbb{R}}}$ such that $(x,H)\in{\mathscr{N}}^{0}$ for some $x\in(u|_{\mathfrak{F}})^{-1}(S)$.
We start with the following classical result.
Lemma 4.7.
  1. (i)
    There exists a finite set $\Omega_{V}=\{V_{1},\ldots,V_{n}\}$ of subspaces of $V_{{\mathbb{R}}}$ with the following property: Each subspace of $V_{{\mathbb{R}}}$ equals $g_{V}V_{j}$ for some $g_{V}\in{\mathrm{GL}}(V_{{\mathbb{R}}})$ and some $j\in\{1,\ldots,n\}$.
  2. (ii)
    There exists a finite set $\Omega_{0}=\{G_{1},\ldots,G_{n}\}$ of semi-simple subgroups of $\mathbf{G}_{0,{\mathbb{R}}}$, with no compact factors, such that the following holds: Each semi-simple subgroups of $\mathbf{G}_{0,{\mathbb{R}}}$ equals $g_{0}G_{j}g_{0}^{-1}$ for some $g_{0}\in\mathbf{G}_{0}({\mathbb{R}})$ and some $j\in\{1,\ldots,n\}$.
Consider the set $\Upsilon$ consisting of elements $(x,g_{V},g_{0},V_{j},G_{j},v)\in(u|_{\mathfrak{F}})^{-1}(S)\times{\mathrm{GL}}(V_{{\mathbb{R}}})\times\mathbf{G}_{0}({\mathbb{R}})\times\Omega_{V}\times\Omega_{0}\times V({\mathbb{R}})$ satisfying the following properties:
  1. (a)
    $g_{0}G_{j}g_{0}^{-1}$ stabilizes $g_{V}V_{j}$; hence $g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1}$ is a subgroup of $\mathbf{G}_{{\mathbb{R}}}=(V\rtimes\mathbf{G}_{0})_{{\mathbb{R}}}$;
  2. (b)
    $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}({\mathbb{R}})}\left((v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)\right)$.
[…]
For each $(x,g_{V},g_{0},V_{j},G_{j},v)\in\Upsilon$, denote by $H^{(j)}_{v}:=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$. Then we obtain a map that is surjective by Lemma 4.7
(4.3) $\psi\colon\Upsilon\rightarrow{\mathscr{N}},\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\left(x,H^{(j)}_{v}\right).$
Composing this map with the functions $d\colon{\mathscr{N}}\rightarrow{\mathbb{R}}$ and $\delta\colon{\mathscr{N}}\rightarrow{\mathbb{R}}$ from (4.1), we obtain two functions on $\Upsilon$
$d_{\Upsilon}\colon\Upsilon\rightarrow{\mathbb{R}},$ $\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\dim_{x}\left((u|_{\mathfrak{F}})^{-1}(S)\bigcap H^{(j)}_{v}({\mathbb{R}})x\right),$
$\delta_{\Upsilon}\colon\Upsilon\rightarrow{\mathbb{R}},$ $\quad(x,g_{V},g_{0},V_{j},G_{j},v)\mapsto\dim_{x}H^{(j)}_{v}({\mathbb{R}})x-\dim_{x}\left((u|_{\mathfrak{F}})^{-1}(S)\bigcap H^{(j)}_{v}({\mathbb{R}})x\right).$
The general theory of o-minimal geometry says that $x\mapsto\dim_{x}X$ is a definable function for any definable set $X$. Hence $d_{\Upsilon}$ and $\delta_{\Upsilon}$ are definable functions on $\Upsilon$.
For the subset ${\mathscr{N}}^{0}\subseteq{\mathscr{N}}$, the inverse $\psi^{-1}({\mathscr{N}}^{0})\subseteq\Upsilon$ is
$\Xi=\big{\{}\normalsize(x,g_{V}$ $,g_{0},V_{j},G_{j},v)\in\Upsilon:\text{ for any }(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime})\in\Upsilon,$
$H^{(j^{\prime})}_{v^{\prime}}({\mathbb{R}})x\subsetneq H^{(j)}_{v}({\mathbb{R}})x\Rightarrow d_{\Upsilon}(x,g_{V},g_{0},V_{j},G_{j},v)>d_{\Upsilon}(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime}),$
$H^{(j)}_{v}({\mathbb{R}})x\subsetneq H^{(j^{\prime})}_{v^{\prime}}({\mathbb{R}})x\Rightarrow\delta_{\Upsilon}(x,g_{V},g_{0},V_{j},G_{j},v)<\delta_{\Upsilon}(x,g^{\prime}_{V},g^{\prime}_{0},V_{j^{\prime}},G_{j^{\prime}},v^{\prime})\big{\}},$
which is a definable subset of $\Upsilon$ since both $d_{\Upsilon}$ and $\delta_{\Upsilon}$ are definable functions.
With these preparations, we are ready to prove Proposition 4.6.
Proof of Proposition 4.6.
Consider the map
$\begin{array}[]{rcl}\rho\colon\Xi&\rightarrow&\bigcup_{i=1}^{n}\left({\mathrm{GL}}(V_{{\mathbb{R}}})/\mathrm{Stab}_{{\mathrm{GL}}(V_{{\mathbb{R}}})}(V_{j})\right)\times\left(\mathbf{G}_{0}({\mathbb{R}})/N_{\mathbf{G}_{0}({\mathbb{R}})}(G_{j})\right)\times V({\mathbb{R}})\\ (x,g_{V},g_{0},V_{j},G_{j},v)&\mapsto&\qquad\qquad(g_{V}V_{j},g_{0}G_{j}g_{0}^{-1},v).\end{array}$
By the surjectivity of $\psi$, for any $(x,H)\in{\mathscr{N}}^{0}$, the group $H$ equals $H^{(j)}_{v}=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$ for some $(g_{V}V_{j},g_{0}G_{j}g_{0}^{-1},v)\in\rho(\Xi)$.
So, it suffices to prove that $\rho(\Xi)$ is finite.
First, the map $\rho$ is clearly definable. Hence $\rho(\Xi)$ is definable.
Next, Proposition 4.5 says that $H^{(j)}_{v}({\mathbb{R}})x$ is a weak Mumford–Tate domain. Hence there exists a ${\mathbb{Q}}$-subgroup $N$ of $\mathbf{G}$ such that $N_{{\mathbb{R}}}=(v,1)(g_{V}V_{j}\rtimes g_{0}G_{j}g_{0}^{-1})(-v,1)$. This implies that $g_{V}V_{j}=(V\cap N)_{{\mathbb{R}}}$ and $g_{0}G_{j}g_{0}^{-1}=N_{0,{\mathbb{R}}}$. Moreover, $(v,1)\left((V\cap N)\rtimes N_{0}\right)(-v,1)=N$ and hence $v\in V({\mathbb{Q}})$. So $\rho(\Xi)$ is countable since ${\mathbb{Q}}$ is countable.
Therefore, $\rho(\Xi)$ is finite because it is definable and countable. ∎
4.4. Proof of Proposition 3.8
Let $Y$ be a weakly optimal subvariety of $S$. Take a complex analytic irreducible component $\tilde{Y}$ of $u^{-1}(Y)$ such that $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. Then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$ by Corollary 4.3. Then by Proposition 4.5, we have $\tilde{Y}^{\mathrm{ws}}=H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}^{0}$.
Hence by the finiteness result Proposition 4.6, there exist finitely many ${\mathbb{Q}}$-groups $N_{1},\ldots,N_{k}<\mathbf{G}$ satisfying the following property: For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}=u(\tilde{Y}^{\mathrm{ws}})$ equals $u(N_{j}({\mathbb{R}})^{+}x)$ for some $j\in\{1,\ldots,k\}$ and some $x\in{\mathscr{D}}$ with $h_{x}({\mathbb{S}})\subseteq\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})$. Then ${\mathscr{D}}^{\prime}_{j}:=\mathrm{N}_{\mathbf{G}}(N_{j})({\mathbb{R}})^{+}x$ is a Mumford–Tate domain and is independent of the choice of such $x$. We are done. ∎
5. Application to non-degeneracy in case of irreducible VHSs
In this section, we give two applications of Theorem 3.1: when the VHS is irreducible and when it has a simple algebraic monodromy group. Both cases apply to the Gross-Shoen and the Ceresa normal functions.
Let $S$ be a smooth irreducible quasi-projective variety. Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. Let $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ be an admissible normal function. Then we have the associated Betti form $\beta_{\nu}$ defined in Definition D.5; it is a semi-positive $(1,1)$-form on $S$.
We wish to check whether $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$, which by Corollary D.7 becomes $r(\nu)=\dim S$ for the Betti rank $r(\nu)$ defined in (1.4). Now Theorem 3.1 gives a checkable criterion. Recall the period map $\varphi=\varphi_{\nu}\colon S\rightarrow\Gamma\backslash{\mathscr{D}}$, the ${\mathbb{Q}}$-group $\mathbf{G}$ and its unipotent radical $V$ which is a vector group from (2.1). We emphasize that $V$ is in general not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$ and its geometric meaning is given by Remark 2.2: $\frac{1}{2}\dim V$ is the relative dimension of ${\mathscr{J}}_{\nu}$, the smallest intermediate Jacobian of sub-VHS of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ translated by a torsion multisection which contains $\nu(S)$.
Notice that the trivial upper bound (2.9) yields the following necessary condition for $r(\nu)=\dim S$:
${\tt(Hyp):}\qquad\dim\varphi(S)=\dim S\leq\frac{1}{2}\dim_{{\mathbb{Q}}}V.$
Theorem 5.1.
Assume: (i) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is irreducible, i.e. the only sub-VHSs are trivial or itself; (ii) $\nu(S)$ is not a torsion section. Then
$r(\nu)=\min\left\{\dim\varphi(S),\frac{1}{2}\dim{\mathbb{V}}_{{\mathbb{Q}},s}\right\}$
for one (and hence for all) $s\in S({\mathbb{C}})$.
In particular, if futhermore $\dim\varphi(S)=\dim S$ and $\dim{\mathbb{V}}_{{\mathbb{Q}},s}\geq 2\dim S$, then we have $\beta_{\nu}^{\wedge\dim S}\not\equiv 0$.
Proof.
Set $\mathbf{G}_{0}=\mathbf{G}/V$. Then $V$ is a $\mathbf{G}_{0}$-submodule of ${\mathbb{V}}_{{\mathbb{Q}},s}$ for one (and hence all) $s\in S({\mathbb{C}})$.
By (i), ${\mathbb{V}}_{{\mathbb{Q}},s}$ is irreducible as a $\mathbf{G}_{0}$-module. By (ii), $V\not=\{0\}$. Hence $V={\mathbb{V}}_{{\mathbb{Q}},s}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. Then $V\cap N$ is a $\mathbf{G}_{0}$-submodule of $V$, and hence gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$. Hence by (i), either $V\cap N=\{0\}$ or $V\cap N=V$.
Assume $V\cap N=\{0\}$. Then $N\lhd\mathbf{G}_{0}=\mathbf{G}/V$ is reductive and it acts trivially on $V=V/(V\cap N)$. Hence, $N({\mathbb{R}})^{+}x_{0}$ is a point for any $x_{0}\in{\mathscr{D}}_{0}$, and therefore $N$ is contained in the center of $\mathbf{G}_{0}$. Using again the fact that $N$ acts trivially on $V$, we see that $N$ is contained in the center of $\mathbf{G}$. So $\dim\varphi_{/N}(S)=\dim\varphi(S)$ in this case.
On the other hand, it is clearly true that $\min_{N,~{}V\cap N=V}\left\{\dim\varphi_{/N}(S)\right\}+\frac{1}{2}\dim V$ is attained at $N=\mathbf{G}$ and hence equals $\frac{1}{2}\dim V$. Now we are done. ∎
Theorem 5.2.
Assume: (i) the connected algebraic monodromy group $H$ of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ is simple; (ii) $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$ has no isotrivial sub-VHS, i.e. locally constant VHS. Then
$r(\nu)=\min\left\{\dim\varphi(S),\frac{1}{2}\dim V\right\}.$
Proof.
Set $\mathbf{G}_{0}=\mathbf{G}/V$. By Deligne, $H\lhd\mathbf{G}^{\mathrm{der}}_{0}$.
Let $N$ be a normal subgroup of $\mathbf{G}$. We may and do assume $N\lhd\mathbf{G}^{\mathrm{der}}$. The reductive part $\mathbf{G}_{N}:=N/(V\cap N)$ is a normal subgroup of $\mathbf{G}$. Now that $\mathbf{G}_{N}\cap H$ is a normal subgroup of $H$, we have either $\mathbf{G}_{N}\cap H=\{1\}$ or $H<\mathbf{G}_{N}$ by (i).
Assume $H<\mathbf{G}_{N}$. Since $\mathbf{G}_{0}$ is reductive, we can decompose $V=(V\cap N)\bigoplus(V\cap N)^{\perp}$ as $\mathbf{G}_{0}$-modules. Now $(V\cap N)^{\perp}$ gives rise to a sub-VHS of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$. But $H$ acts trivially on $(V\cap N)^{\perp}$ because $\mathbf{G}_{N}$ acts trivially on $V/(V\cap N)$. So $(V\cap N)^{\perp}=\{0\}$ by (ii). Hence $V\cap N=V$, and hence $V<N$. Thus $\varphi_{/N}(S)$ is a point. So $\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)=\frac{1}{2}\dim V$.
If $\mathbf{G}_{N}\cap H=\{1\}$, then $\dim\varphi_{/N}(S)=\dim\varphi(S)$. Now we are done. ∎
Now, we apply these two theorems to the Gross-Shoen and the Ceresa normal functions. Let $\nu_{{\mathrm{GS}}}$ (resp. $\nu_{{\mathrm{Ce}}}$) be the Gross–Schoen (resp. Ceresa) normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.5), with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10. Let $\beta_{{\mathrm{GS}}}$ and $\beta_{{\mathrm{Ce}}}$ be the associated Betti forms.
Corollary 5.3.
Assume $g\geq 3$. The Betti ranks $r(\nu_{{\mathrm{GS}}})=r(\nu_{{\mathrm{Ce}}})=3g-3$. Equivalently, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$ and $\beta_{{\mathrm{Ce}}}^{\wedge(3g-3)}\not\equiv 0$.
Proof.
We shall apply Theorem 5.1 and let us check the assumptions. Assumption (i) holds by Proposition C.10. (iii). Assumption (ii) holds by Proposition C.10. (ii). Moreover the VHS ${\mathbb{V}}_{{\mathbb{Z}}}$ on $S$ has maximal moduli, so $\dim\varphi_{{\mathrm{GS}}}({\mathscr{M}}_{g})=\dim\varphi_{{\mathrm{Ce}}}({\mathscr{M}}_{g})=\dim{\mathscr{M}}_{g}=3g-3$ where $\varphi_{{\mathrm{GS}}}$ and $\varphi_{{\mathrm{Ce}}}$ are the period maps. Finally, $\dim{\mathbb{V}}_{{\mathbb{Q}},s}=\frac{2g(2g-1)(2g-2)}{6}-2g$, which is $>2(3g-3)$ when $g\geq 3$. Hence, we can conclude by Theorem 5.1. ∎
We are ready to prove Theorem 1.3.
Proof of Theorem 1.3.
Denote for simplicity by $S={\mathscr{M}}_{g}$. Let $\nu$ be either $\nu_{{\mathrm{GS}}}$ or $\nu_{{\mathrm{Ce}}}$.
Set $U:=S\setminus S_{{\mathscr{F}}}(1)$, which is Zariski open by Theorem 3.2.
By Corollary 5.3 and the “In particular” part of Theorem 3.2, we have that $\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}=0$. Thus $S_{{\mathscr{F}}}(1)\not=S$. Now, part (i) follows from Proposition C.7.
For (ii), notice that by definition $S_{{\mathscr{F}}}(1)$ contains any analytic curve $C\subseteq S^{\mathrm{an}}$ such that $\nu(C)$ is torsion. Thus, the set $\{s\in U({\mathbb{C}}):\nu(s)\text{ is torsion}\}$ is discrete, and hence at most countable. ∎
6. Northcott property for Gross–Schoen and Ceresa cycles
In this section, we turn to the arithmetic applications.
6.1. Construction of the adelic line bundle
Let ${\mathscr{M}}_{g}$ be the moduli scheme of smooth curves of genus $g$ over ${\mathbb{Z}}$, and let ${\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve.
Denote by ${\mathscr{J}}_{g}=\mathrm{Jac}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ the relative Jacobian. Identify ${\mathscr{J}}_{g}$ with its dual via the principal polarization given by a suitable theta divisor.
The Poincaré line bundle ${\mathscr{P}}$ on ${\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$ extends to an integrable adelic line bundle $\overline{{\mathscr{P}}}$ as follows. Define ${\mathscr{P}}^{\Delta}:=\Delta^{*}{\mathscr{P}}$ for the diagonal $\Delta\colon{\mathscr{J}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$. Then ${\mathscr{P}}^{\Delta}$ is relatively ample on ${\mathscr{J}}_{g}\rightarrow{\mathscr{M}}_{g}$, and $[2]^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{\otimes 4}$. So $({\mathscr{J}}_{g},[2],{\mathscr{P}}^{\Delta})$ is a polarized dynamical system over ${\mathscr{M}}_{g}$ in the sense of [YZ21, §$\S$2.6.1]. Thus, Tate’s limit process gives a nef adelic line bundle $\overline{{\mathscr{P}}}^{\Delta}$ on ${\mathscr{J}}_{g}$, as executed by [YZ21, Thm. 6.1.1]. Now we obtain the desired $\overline{{\mathscr{P}}}\in\widehat{\mathrm{Pic}}({\mathscr{J}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}$ by letting
$2\overline{{\mathscr{P}}}:=m^{*}\overline{{\mathscr{P}}}^{\Delta}-p_{1}^{*}\overline{{\mathscr{P}}}^{\Delta}-p_{2}^{*}\overline{{\mathscr{P}}}^{\Delta},$
where $m,p_{1},p_{2}\colon{\mathscr{J}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$ with $m$ being the addition and $p_{1}$ (resp. $p_{2}$) being the projection to the first (resp. second) factor.
Take $\mathbf{\xi}\in\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ such that $(2g-2)\mathbf{\xi}=\omega_{{\mathscr{C}}_{g}/{\mathscr{M}}_{g}}$. Let
$i_{\mathbf{\xi}}\colon{\mathscr{C}}_{g}\longrightarrow{\mathscr{J}}_{g}$
be the Abel–Jacobi map based at $\mathbf{\xi}$. Then we have an ${\mathscr{M}}_{g}$-morphism $(i_{\mathbf{\xi}},i_{\mathbf{\xi}})\colon{\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}\rightarrow{\mathscr{J}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{J}}_{g}$, and hence get an integrable adelic line bundle on ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}$
$\overline{{\mathscr{Q}}}:=(i_{\mathbf{\xi}},i_{\mathbf{\xi}})^{*}\overline{{\mathscr{P}}}\in\widehat{\mathrm{Pic}}({\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}},$
and we can compute
(6.1) ${\mathscr{Q}}={\mathscr{O}}(\Delta)-p_{1}^{*}\xi-p_{2}^{*}\xi.$
with $\Delta$ the diagonal of ${\mathscr{C}}_{g}\rightarrow{\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}$, and $p_{1},p_{2}$ the projections ${\mathscr{C}}_{g}\times_{{\mathscr{M}}_{g}}{\mathscr{C}}_{g}\rightarrow{\mathscr{C}}_{g}$.
Finally the Deligne pairing gives an adelic line bundle on ${\mathscr{M}}_{g}$ by [YZ21, Thm. 4.1.3]
(6.2) ¯$\overline{{\mathscr{L}}}:=\langle\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}},\overline{{\mathscr{Q}}}\rangle\in\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}.$
The following theorem is a reformulation of the second-named author’s [Zha10, Thm. 2.3.5].
Theorem 6.1.
For any $s\in{\mathscr{M}}_{g}(\overline{\mathbb{Q}})$, we have
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}=h_{\overline{{\mathscr{L}}}}(s).$
Proof.
In some neighborhood $U$ of $s$ in ${\mathscr{M}}_{g}$, $\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})=\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is an element of $Z^{2}({\mathscr{C}}_{U}^{3})$, i.e. it is a $2$-cocycle of ${\mathscr{C}}_{U}^{3}$.
Let $\ell$ be a rational section of ${\mathscr{Q}}$ over ${\mathscr{C}}_{U}^{2}$. By (6.1), the divisor $\mathrm{div}(\ell_{s^{\prime}})$ at each $s^{\prime}\in U({\mathbb{C}})$ is not the pullback of a divisor under the two natural projections ${\mathscr{C}}_{U}^{2}\rightarrow{\mathscr{C}}_{U}$. So $\mathrm{div}(\ell)$ can be seen as a correspondence of ${\mathscr{C}}_{U}$. Up to shrinking $U$ we can take rational sections $\ell_{1},\ell_{2},\ell_{3}$ of ${\mathscr{Q}}$ over ${\mathscr{C}}_{U}^{2}$ with the following property: For their divisors $t_{1},t_{2},t_{3}\in\mathrm{Div}({\mathscr{C}}_{U}^{2})$, we have
$|t_{1}|\cap|t_{2}|\cap|t_{3}|=\emptyset,\qquad|\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|\cap|(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})|=\emptyset,$
where $t_{1}\otimes t_{2}\otimes t_{3}$ is seen as a correspondence of ${\mathscr{C}}_{U}^{3}$. Notice that we have a rational section $\langle\ell_{1},\ell_{2},\ell_{3}\rangle$ of ${\mathscr{L}}$ on $U$.
We shall apply [Zha10, Thm. 2.3.5] to ${\mathscr{C}}_{s}$. We have seen that $t_{1},t_{2},t_{3}$ restricted to the fiber over $s\in U(\overline{\mathbb{Q}})$ satisfies the assumption of [Zha10, Thm. 2.3.5], so
(6.3) $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),(t_{1,s}\otimes t_{2,s}\otimes t_{3,s})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}=\hat{t}_{1,s}\cdot\hat{t}_{2,s}\cdot\hat{t}_{3,s}$
where $\hat{t}_{1,s}$, $\hat{t}_{2,s}$, $\hat{t}_{3,s}$ are suitable arithmetic divisors on some model of ${\mathscr{C}}_{s}^{2}$ extending $t_{1,s}$, $t_{2,s}$, $t_{3,s}$. The right hand side is precisely $\frac{1}{[{\mathbb{Q}}(s):{\mathbb{Q}}]}\widehat{\deg}\overline{{\mathscr{L}}}|_{s}=h_{\overline{{\mathscr{L}}}}(s)$. The left hand side is $\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s}),\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})\rangle_{\mathrm{BB}}$ because, by (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$. Hence, we are done. ∎
6.2. Relating with the Gross–Schoen normal function
Let $\nu_{{\mathrm{GS}}}$ be the Gross–Schoen normal function ${\mathscr{M}}_{g}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in (C.5) (with ${\mathbb{V}}_{{\mathbb{Z}}}$ from Proposition C.10).
By (6.1), $(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ is rationally equivalent to $\Delta_{{\mathrm{GS}},\mathbf{\xi}}({\mathscr{C}}_{U})$ Hence, they define the same normal functions $\nu_{{\mathrm{GS}}}$ on $U$. Set
$\overline{{\mathscr{P}}^{{\mathrm{GS}}}}:=(\nu_{{\mathrm{GS}}},\nu_{{\mathrm{GS}}})^{*}\overline{{\mathscr{P}}}=\nu_{{\mathrm{GS}}}^{*}\overline{{\mathscr{P}}^{\Delta}}$
for the metrized Poincaré bundle $\overline{{\mathscr{P}}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{{\mathscr{M}}_{g}}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defined in Definition D.1 (or the metrized tautological bundle $\overline{{\mathscr{P}}^{\Delta}}$ from Definition D.3). Set
$\beta_{{\mathrm{GS}}}:=c_{1}(\overline{{\mathscr{P}}^{{\mathrm{GS}}}})$
to be the Betti form, i.e. the curvature of $\overline{{\mathscr{P}}^{{\mathrm{GS}}}}$. It is a semi-positive $(1,1)$-form. The following Proposition can also be deduced from R. de Jong’s work [dJ16, (9.3)]:
Proposition 6.2.
We have the following identity of $(1,1)$-forms on ${\mathscr{M}}_{g}$:
$c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})=\beta_{{\mathrm{GS}}}.$
In particular, $c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})$ is semi-positive.
Proof.
We use the notation in the proof of Theorem 6.1. For any $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, we obtain a neighborhood $U$ such that (6.3) holds for any $s^{\prime}\in U({\mathbb{C}})$. So we get
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U}),(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})\rangle_{\infty}=-\log\|\langle\ell_{1},\ell_{2},\ell_{3}\rangle\|_{\infty}\quad\text{ on }U.$
By Proposition D.2, we have a section $\beta\in H^{0}(U,{\mathscr{P}}^{{\mathrm{GS}}})$ such that
$\langle\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U}),(t_{1}\otimes t_{2}\otimes t_{3})^{*}\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{U})\rangle_{\infty}=-\log\|\beta\|\quad\text{ on }U.$
So $\log\|\beta\|=\log\|\langle\ell_{1},\ell_{2},\ell_{3}\rangle\|_{\infty}$ on $U$. Taking ¯$\frac{\partial\bar{\partial}}{\pi i}$, we get that $c_{1}(\overline{{\mathscr{P}}^{{\mathrm{GS}}}})=c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})$ on $U$. Now we can conclude by letting $s$ run over ${\mathscr{M}}_{g}({\mathbb{C}})$. ∎
6.3. Bigness of the generic fiber of the adelic line bundle
By [YZ21, Thm. 5.3.5], lower bounds of the height function $h_{\overline{{\mathscr{L}}}}$ correspond to bigness properties of $\overline{{\mathscr{L}}}$. In this paper, we prove the bigness of the generic fiber of $\overline{{\mathscr{L}}}$ and deduce the desired height comparison from it.
There is a natural base change of adelic line bundles $\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Z}})_{{\mathbb{Q}}}\rightarrow\widehat{\mathrm{Pic}}({\mathscr{M}}_{g}/{\mathbb{Q}})_{{\mathbb{Q}}}$. Denote by $\tilde{{\mathscr{L}}}$ the image of $\overline{{\mathscr{L}}}$ under this base change map; $\tilde{{\mathscr{L}}}$ is called the generic fiber of $\overline{{\mathscr{L}}}$.
We start by stating the following formula for $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})$. Notice that it does not follow from arithmetic Hilbert–Samuel because we have not proved the nefness of $\tilde{{\mathscr{L}}}$.
Proposition 6.3.
We have
$\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}c_{1}(\overline{{\mathscr{L}}}_{{\mathbb{C}}})^{3g-3}.$
The proof of this proposition shall be postponed to §$\S$6.4.
Theorem 6.4.
Assume $g\geq 3$. Then the adelic line bundle $\tilde{{\mathscr{L}}}$ is big, i.e. $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$.
Proof.
By Proposition 6.3 and Proposition 6.2, we have $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})=\int_{{\mathscr{M}}_{g}({\mathbb{C}})}\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}$. Now $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\geq 0$ since $\beta_{{\mathrm{GS}}}$ is semi-positive. By Corollary 5.3, $\beta_{{\mathrm{GS}}}^{\wedge(3g-3)}\not\equiv 0$. Hence $\widehat{\mathrm{vol}}(\tilde{{\mathscr{L}}})>0$. ∎
Proof of Theorem 1.1.
The claim for $\langle{{\mathrm{GS}}({\mathscr{C}}_{s}),{\mathrm{GS}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ is a direct consequence of Theorem 6.4 and [YZ21, Thm. 5.3.5.(iii)]. Notice that one can either take the adelic line bundle on ${\mathscr{M}}_{g}$, which defines the Faltings height (it exists by [YZ21, §$\S$2.6.2]), or one can take any ample line bundle on a suitable compactification of ${\mathscr{M}}_{g}$ and then use the comparison of the logarithmic Weil height and the Faltings height.
The claim for $\langle{{\mathrm{Ce}}({\mathscr{C}}_{s}),{\mathrm{Ce}}({\mathscr{C}}_{s})}\rangle_{\mathrm{BB}}$ follows because, by work of the second-named author [Zha10], the Ceresa cycles and Gross–Schoen have the same height up to some positive multiple. ∎
6.4. Proof of Proposition 6.3
Denote for simplicity by ${\mathscr{S}}={\mathscr{M}}_{g}$.
By the flatness of extension ${\mathbb{Q}}\subseteq{\mathbb{C}}$, we have $\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{Q}})=\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{\mathbb{C}})$. By [YZ21, Theorem 5.2.1], it suffices to construct a sequence of model Hermitian line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i})$ with limit $({\mathscr{S}},\overline{\mathscr{L}})$ so that
$\lim_{i{\longrightarrow}\infty}\widehat{\mathrm{vol}}(\tilde{\mathscr{L}}_{i})=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})^{3g-3}.$
Let $\overline{\mathscr{N}}$ be an ample Hermitian line bundle on some integral model ${\mathscr{S}}$ of ${\mathscr{S}}$. Following proof of [YZ21, Theorem 6.1.1], we have a sequence of model line bundles $({\mathscr{J}}_{i},\overline{\mathscr{P}}^{\Delta}_{i},\ell_{i})$ with the limit $({\mathscr{J}},\overline{\mathscr{P}}^{\Delta})$, so that $\overline{\mathscr{P}}^{\Delta}_{i}+4^{-i}\pi^{*}\overline{\mathscr{N}}$ is nef.
For any $n\in{\mathbb{N}}$, we have an action of ${\mathscr{J}}[n]$ on ${\mathscr{J}}$ by translating torsion points:
$m_{n}\colon{\mathscr{J}}[n]\times{\mathscr{J}}\overset{(m,p)}{{\longrightarrow}}{\mathscr{J}}\times{\mathscr{J}},\qquad(t,x)\mapsto(x+t,x)$
By the Theorem of square, ${\mathrm{N}}_{p}m_{n}^{*}{\mathscr{P}}^{\Delta}=({\mathscr{P}}^{\Delta})^{n^{2g}}$. Thus, we obtain an adelic metrized line bundle
$\overline{\mathscr{P}}^{\Delta}_{i,n}:=n^{-2g}{\mathrm{N}}_{p}m_{n}^{*}\overline{\mathscr{P}}^{\Delta}_{i}.$
This bundle is again realized on some model ${\mathscr{J}}_{i,n}$ of ${\mathscr{J}}$ with connection morphism $\ell_{i,n}:{\mathscr{P}}^{\Delta}{\longrightarrow}{\mathscr{P}}^{\Delta}_{i,n,{\mathscr{J}}}$. Moreover ${\mathrm{div}}(\ell_{i,n})$ is in fact bounded by ${\mathrm{div}}(\ell_{i})$. Over ${\mathbb{C}}$, the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,n,{\mathbb{C}}})$ is obtained from the curvature form $c_{1}(\overline{\mathscr{P}}^{\Delta}_{i,{\mathbb{C}}})$ by taking over average over $n$-torsion points. It follows that these forms converge to $c_{1}(\overline{\mathscr{P}}^{\Delta})$ uniformly in any compact subset of ${\mathscr{J}}({\mathbb{C}})$. These bundles also induce a double sequence of model line bundles $({\mathscr{S}}_{i},\overline{\mathscr{L}}_{i,n})$ of $({\mathscr{S}},\overline{\mathscr{L}})$ so that they convergent to $({\mathscr{S}},\overline{\mathscr{L}})$ as $i{\longrightarrow}\infty$, and that the metric $c_{1}(\overline{\mathscr{L}}_{i,n,{\mathbb{C}}})$ uniformly convergent to $c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})$ as $n{\longrightarrow}\infty$.
More precisely, we let $\Omega_{i}$ be an increasing sequence of relatively compact open subsets of ${\mathscr{S}}({\mathbb{C}})$ so that ${\mathscr{S}}({\mathbb{C}})=\cup\Omega_{i}$ and $\epsilon_{i}$ be a decreasing sequence of positive numbers convergent to 00 so that on $\Omega_{i}$,
$c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{-1}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Then for each $i$, we choose $n_{i}$ so that
(6.4) $-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i,n_{i},{\mathbb{C}}})-c_{1}(\overline{\mathscr{L}}_{\mathbb{C}})\leq\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$
as hermitian forms on the tangent bundle on $\Omega_{i}$. We simply write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{L}}_{i,n_{i}}$. Using the reference curvature $c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})$, we may talk about eigenvalues of $c_{1}(\overline{\mathscr{L}}_{{\mathbb{C}}})$ and $c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})$.
Now we apply Demailly’s Morse inequality [Dem91] for the bundle ¯$\overline{\mathscr{L}}_{i}$ on ${\mathscr{S}}_{i}({\mathbb{C}})$. For each $q\in{\mathbb{N}}$, let ${\mathscr{S}}_{i,q}$ denote the subset of ${\mathscr{S}}_{i}({\mathbb{C}})$ of points where $c_{1}(\overline{\mathscr{L}}_{i})$ has $q$-negative eigenvalues and $n-q$ positive eigenvalues. Then by [Dem91, (1.3), (1.5)], we have the following estimate as $k{\longrightarrow}\infty$:
$h^{q}(k{\mathscr{L}}_{i})\leq\frac{k^{d}}{d!}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|+o(k^{d}),\qquad d:=3g-3$
$\sum_{q}(-1)^{q}h^{q}(k{\mathscr{L}}_{i})=\frac{k^{d}}{d!}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}+o(k^{d}).$
It follows that
$\left|\widehat{\mathrm{vol}}({\mathscr{L}}_{i})-\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}\right|\leq\sum_{q>0}\left|\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}\right|.$
By [YZ21, Theorem 5.4.4],
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}}_{i})^{d}=\int_{{\mathscr{S}}({\mathbb{C}})}c_{1}({\mathscr{L}})^{d}.$
Thus, it remains to the following estimate for each $q>0$:
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}_{i,q}}c_{1}(\overline{\mathscr{L}}_{i})^{d}=0.$
We want to prove this on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$ and on its complement ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$ respectively.
By (6.4), we have
$-\epsilon_{i}^{d}c_{1}(\overline{\mathscr{N}}_{\mathbb{C}})\leq c_{1}(\overline{\mathscr{L}}_{i})\leq(\epsilon^{-1}+\epsilon^{d})c_{1}(\overline{\mathscr{N}}_{\mathbb{C}}).$
Thus on $\Omega_{i}\cap{\mathscr{S}}_{i,q}$, $c_{1}(\overline{\mathscr{L}}_{i})$ has all eigenvalue $\leq\epsilon^{-1}+\epsilon^{d}$ and one negative eigenvalue with absolute value bounded by $\leq\epsilon^{d}$. It follows that $|c_{1}({\mathscr{L}}_{i})^{d}|$ is bounded by
$\epsilon^{d}(\epsilon^{-1}+\epsilon^{d})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}=\epsilon_{i}(1+\epsilon_{i}^{d+1})^{d-1}c_{1}(\overline{\mathscr{N}})^{d}.$
It follows that
$\int_{{\mathscr{S}}_{i,q}\cap\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|=O(\epsilon_{i}).$
It remains to treat the integral over ${\mathscr{S}}_{i,q}\setminus\Omega_{i}$. Using decomposition
${\mathscr{P}}=\frac{1}{2}(m^{*}{\mathscr{P}}^{\Delta}-p_{1}^{*}{\mathscr{P}}^{\Delta}-p_{2}^{*}{\mathscr{P}}^{\Delta}),$
we may write ¯$\overline{\mathscr{L}}_{i}=\overline{\mathscr{E}}_{i}-\overline{\mathscr{F}}_{i}$, where ¯$\overline{\mathscr{E}}_{i}$ and ¯$\overline{\mathscr{F}}_{i}$ are two sequences of new bundles convergent to $\overline{\mathscr{E}}$ and $\overline{\mathscr{F}}$ with smooth metrics respectively. Then we have
$\int_{{\mathscr{S}}_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
Now let $\Omega_{i}^{\prime}\subset\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}={\mathscr{S}}({\mathbb{C}})$. Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}(x)=1$ on ${\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}=0$ on $\Omega_{i}^{\prime}$. Then for any $i\geq j$ we have
$\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
Fix $j$ and take $i{\longrightarrow}\infty$, we get
$\limsup_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}.$
Let $j{\longrightarrow}\infty$. We get
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}=0.$
We are done. ∎
Appendix A Variation of mixed Hodge structures
A.1. Definitions
Let $R$ be a subring of ${\mathbb{R}}$.
Definition A.1.
Let $M$ be a free $R$-module of finite rank.
  1. (i)
    An $R$-pure Hodge structure on $M$ of weight $n$ is a decreasing filtration $F^{\bullet}$ on $M_{{\mathbb{C}}}$ (the Hodge filtration) such that $M_{{\mathbb{C}}}=F^{p}M_{{\mathbb{C}}}\oplus\overline{F^{n+1-p}M_{{\mathbb{C}}}}$ for all $p\in{\mathbb{Z}}$.
  2. (ii)
    An $R$-mixed Hodge structure on $M$ is a triple $(M,W_{\bullet},F^{\bullet})$ consisting of two filtrations, an increasing filtration $W_{\bullet}$ on $M_{{\mathbb{Q}}}$ (the weight filtration) and a decreasing filtration $F^{\bullet}$ on $M_{{\mathbb{C}}}$ (the Hodge filtration), such that for each $k\in{\mathbb{Z}}$, $\mathrm{Gr}_{k}^{W}M_{{\mathbb{Q}}}=W_{k}/W_{k-1}$ is a ${\mathbb{Q}}$-pure Hodge structure of weight $k$ for the filtration on $\mathrm{Gr}_{k}^{W}(M_{\mathbb{C}})$ deduced from $F^{\bullet}$.
Pure Hodge structures of weight $n$ can be defined in terms of bigradings. Indeed, set $M^{p,n-p}:=F^{p}M_{{\mathbb{C}}}\cap\overline{F^{p+1}M_{{\mathbb{C}}}}$, then $M_{{\mathbb{C}}}=\oplus_{p}M^{p,n-p}$ (the Hodge decomposition) and $\overline{M^{n-p,p}}=M^{p,n-p}$. We have $F^{p}=\oplus_{p^{\prime}\geq p}M^{p^{\prime},n-p^{\prime}}$.
For a mixed Hodge structure $(M,W_{\bullet},F^{\bullet})$, the numbers $k\in{\mathbb{Z}}$ such that $\mathrm{Gr}_{k}^{W}M_{{\mathbb{Q}}}\not=0$ are called its weights, and the numbers $h^{p,q}(M)=\dim_{{\mathbb{C}}}F^{p}\mathrm{Gr}^{W}_{p+q}(M_{{\mathbb{C}}})/F^{p+1}\mathrm{Gr}^{W}_{p+q}(M_{{\mathbb{C}}})$ are called its Hodge numbers .
For each $n\in{\mathbb{Z}}$, define $R(n)$ to be the pure Hodge structure on $R$ of weight $-2n$ such that $R(n)^{-n,-n}={\mathbb{C}}$ and $R(n)^{p,q}=0$ for all $(p,q)\not=(-n,-n)$.
A polarization on a pure Hodge structure $V$ of weight $n$ is a morphism of Hodge structures
$Q\colon V_{{\mathbb{Q}}}\otimes V_{{\mathbb{Q}}}\longrightarrow{\mathbb{Q}}(-n)$
such that the Hermitian form on $V_{{\mathbb{C}}}$ given by $Q(Cu,\bar{v})$ is positive-definite where $C$ is the Weil operator ($C|_{H^{p,q}}=i^{p-q}$ for all $p,q$).
A.2. Mumford–Tate group
Now, let us turn to a more group theoretical point of view on mixed Hodge structures. Let ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$ be the Deligne torus, i.e. the real algebraic group such that ${\mathbb{S}}({\mathbb{R}})={\mathbb{C}}^{*}$ and ${\mathbb{S}}({\mathbb{C}})={\mathbb{C}}^{*}\times{\mathbb{C}}^{*}$, and that the complex conjugation on ${\mathbb{S}}({\mathbb{C}})$ sends $(z_{1},z_{2})\mapsto(\bar{z}_{2},\bar{z}_{1})$.
As for pure Hodge structures, mixed Hodge structures can also be equivalently defined in terms of bigradings by Deligne [Del71, 1.2.8]. Given a ${\mathbb{Q}}$-vector space $M$ of finite dimension, a bigrading $M_{{\mathbb{C}}}=\oplus_{p,q\in{\mathbb{Z}}}I^{p,q}$ is equivalent to a homomorphism $h\colon{\mathbb{S}}_{{\mathbb{C}}}\rightarrow\mathrm{GL}(M_{{\mathbb{C}}})$. In particular, any mixed Hodge structure on $M$ defines a unique homomorphism $h\colon{\mathbb{S}}_{{\mathbb{C}}}\rightarrow\mathrm{GL}(M_{{\mathbb{C}}})$, and we use $(M,h)$ to denote this mixed Hodge structure.
Definition A.2.
For any mixed Hodge structure $(M,h)$, its Mumford–Tate group is the smallest ${\mathbb{Q}}$-subgroup $\mathbf{G}$ of $\mathrm{GL}(M_{{\mathbb{Q}}})$ such that $h({\mathbb{S}}_{{\mathbb{C}}})\subseteq\mathbf{G}({\mathbb{C}})$.
Now we assume that $M$ has weight 00 and $-1$ with $M_{-1}:={\mathrm{Gr}}^{W}_{-1}M$ and $M_{0}:={\mathrm{Gr}}^{W}_{0}M$. Then, we have an exact sequence
$00{\longrightarrow}M_{-1}{\longrightarrow}M{\longrightarrow}M_{0}{\longrightarrow}$ 0 .
It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $N$. Let $\mathbf{G}_{0}$ denote the Mumford–Tate group of $M_{-1}$. Then, we have an exact sequence of reductive groups:
$0{\longrightarrow}V{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
where $V$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$. If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne. Thus, $V$ is the unipotent radical of $\mathbf{G}$.
The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=M_{-1,{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$. This induces splittings
$\mathbf{G}_{\mathbb{R}}=V_{\mathbb{R}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
Moreover $h({\mathbb{S}})$ is included into $G_{0,{\mathbb{R}}}$. Thus, we have proved the following:
Lemma A.3.
Assume $(M,h)$ has weight $-1$ and 00. The following holds.
  1. (i)
    The $h$ is defined over ${\mathbb{R}}$.
  2. (ii)
    The unipotent radical $V$ of $\mathbf{G}$ is a vector group.
A.3. Variation of mixed Hodge structures and admissibility
Definition A.4.
Let $S$ be a connected complex manifold. A variation of mixed Hodge structures (VMHS) on $S$ is a triple $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ consisting of
    a local system ${\mathbb{M}}_{{\mathbb{Z}}}$ of free ${\mathbb{Z}}$-modules of finite rank on $S$ ,
    a finite increasing filtration (weight filtration) $W_{\bullet}$ of the local system ${\mathbb{M}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{Q}}_{S}$ by local subsystems,
    a finite decreasing filtration (Hodge filtration) ${\mathscr{F}}^{\bullet}$ of the holomorphic vector bundle ${\mathscr{M}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathscr{O}}_{S}$ by holomorphic subbundles
satisfying the following properties
  1. (i)
    for each $s\in S$, the triple $({\mathbb{M}}_{s},W_{\bullet},{\mathscr{F}}^{\bullet}_{s})$ defines a mixed Hodge structure on ${\mathbb{M}}_{s}$,
  2. (ii)
    the connection $\nabla\colon{\mathscr{M}}\rightarrow{\mathscr{M}}\otimes_{{\mathscr{O}}_{S}}\Omega_{S}^{1}$ whose sheaf of horizontal sections is ${\mathbb{M}}_{{\mathbb{C}}}:={\mathbb{M}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{C}}_{S}$ satisfies the Griffiths’ transversality condition
    $\nabla({\mathscr{F}}^{p})\subseteq{\mathscr{F}}^{p-1}\otimes\Omega_{S}^{1}.$
The weights and Hodge numbers of $({\mathbb{M}}_{s},W_{\bullet},{\mathscr{F}}^{\bullet}_{s})$ are the same for all $s\in S$. We call them the weights and the Hodge numbers of the VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$.
If there is only one $n\in{\mathbb{Z}}$ such that $\mathrm{Gr}^{W}_{n}{\mathbb{M}}\not=0$, then each fiber of this VMHS is a pure Hodge structure of weight $n$. In this case, the VMHS is said to be pure. More precisely, we have the following definition.
Definition A.5.
A variation of Hodge structures (VHS) of weight $n$ on $S$ is a pair $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ such that $({\mathbb{V}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ is a VMHS, where $W_{\bullet}$ is the increasing filtration on ${\mathbb{V}}:={\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}_{S}}{\mathbb{Q}}_{S}$ defined by $W_{n-1}=0$ and $W_{n}={\mathbb{V}}$.
To each VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ on $S$, we can associate variations of pure Hodge structures obtained from $\mathrm{Gr}_{k}^{W}{\mathbb{M}}$.
We shall use the following convention: For each $n\in{\mathbb{Z}}$ and any VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$, define ${\mathbb{V}}_{{\mathbb{Z}}}(n)$ to be the VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet-n})$. In particular ${\mathbb{Z}}(n)_{S}$ be the VHS on $S$ of weight $-2n$ such that $({\mathbb{Z}}(n)_{S})_{s}={\mathbb{Z}}(n)$ for each $s\in S$.
Definition A.6.
A polarization of VHS $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ on $S$ of weight $n$ is a morphism of VHS ${\mathbb{V}}\otimes{\mathbb{V}}\rightarrow{\mathbb{Q}}(-n)_{S}$ inducing on each fiber a polarization.
We say that a VMHS $({\mathbb{M}}_{{\mathbb{Z}}},W_{\bullet},{\mathscr{F}}^{\bullet})$ is graded-polarizable if $\mathrm{Gr}_{k}^{W}{\mathbb{M}}$ has a polarization for each $k\in{\mathbb{Z}}$.
Example A.7.
Let $f\colon X\rightarrow S$ be a projective smooth morphism of algebraic varieties over ${\mathbb{C}}$ with irreducible fibers of dimension $d$. For each $s\in S({\mathbb{C}})$, the cohomology group $H^{n}(X_{s},{\mathbb{Z}})$ is endowed with a natural Hodge structure of weight $n$ by the de Rham–Betti comparison, and this makes $R^{n}f_{*}{\mathbb{Z}}_{X}$ into a VHS on $S^{\mathrm{an}}$ of weight $n$. It is polarizable by the hard Lefschetz theorem and the Lefschetz decomposition.
We close this subsection with a discussion on the admissibility of VMHS. Let $S$ be a smooth, complex quasi-projective variety. Given a VMHS on $S^{\mathrm{an}}$, we often need to extend it to suitable compactifications of $S^{\mathrm{an}}$ and hence study the asymptotic behavior of the VMHS near the boundary. This leads to the definition of admissible VMHS, which are the graded-polarized VMHSs with good asymptotic properties. This concept was introduced by Steenbrick–Zucker [SZ85, Prop. 3.13] on a curve and Kashiwara [Kas86, 1.8 and 1.9] in general, and the property for a VMHS on $S^{\mathrm{an}}$ to be admissible does not depend on the choice of the compactification. We shall not recall the precise definition here (see, for example, [PS08, Defn. 14.49]), but point out the following:
  1. (1)
    Any VMHS arising from geometry is admissible [EZ86] (see also [BZ14]).
  2. (2)
    Any VHS is admissible.
Appendix B Classfying space and Mumford–Tate domain
We relate the Betti foliation on intermediate Jacobians to the fibered structure of certain Mumford–Tate domain parametrizing ${\mathbb{Q}}$-mixed Hodge structure of weight $-1$ and 00. In this appendix, we recall and prove some results about such classifying spaces and Mumford–Tate domains. The main result is Proposition B.3, which explains and compares the semi-algebraic structure and the complex structure of such Mumford–Tate domains. This comparison is used to study the Betti foliation of intermediate Jacobians.
B.1. Classifying space
Let $M_{-1}$ be a finite dimensional ${\mathbb{Q}}$-vector space. Let $M=M_{-1}\oplus{\mathbb{Q}}$.
B.1.1. Pure Hodge structures
Consider the polarized Hodge data on $M_{-1}$: a non-degenerate skew pairing $Q_{-1}\colon M_{-1}\otimes M_{-1}\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{M_{-1}}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{-1,{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{M_{-1}}=h^{q,p}_{M_{-1}}$. Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $M_{-1}$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $M_{-1,{\mathbb{C}}}$ has complex dimension $h^{p,q}$. Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(M_{-1},Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
${\mathscr{M}}_{0}=\mathbf{G}_{0}({\mathbb{R}})^{+}x_{0}$
for any point $x_{0}\in{\mathscr{M}}_{0}$. This makes ${\mathscr{M}}_{0}$ into a semi-algebraic open subset of a flag variety ${\mathscr{M}}_{0}^{\vee}$, which is a suitable $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{C}})$-orbit, and hence endows ${\mathscr{M}}_{0}$ with a complex structure.
We can be more explicit on the action of $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ on ${\mathscr{M}}_{0}$. For each $x_{0}\in{\mathscr{M}}_{0}$, we have a Hodge decomposition and a Hodge filtration $F^{\bullet}_{x_{0}}$
(B.1) $M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(M_{-1,x_{0}})^{p,q},\qquad F^{p}_{x_{0}}M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(M_{-1,x_{0}})^{p^{\prime},q^{\prime}}$
with $(M_{-1,x_{0}})^{q,p}=\overline{(M_{-1,x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(M_{-1,{\mathbb{R}}})$, with $(M_{-1,x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$. It is known that $h_{x_{0}}({\mathbb{S}})<\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})$ for all $x_{0}\in{\mathscr{M}}_{0}$. Hence we have a $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective,
(B.2) ${\mathscr{M}}_{0}\rightarrow{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}),\quad x_{0}\mapsto h_{x_{0}}$
with the action of $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ on ${\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}})$ given by conjugation. So we will view ${\mathscr{M}}_{0}$ as a subset of ${\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}})$.
We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $M_{-1}$ determined by $x_{0}$. We have that $\mathrm{MT}_{x_{0}}$ is a subgroup of $\mathbf{G}_{0}^{{\mathscr{M}}}$ for all $x_{0}\in{\mathscr{M}}_{0}$ and equals $\mathbf{G}_{0}^{{\mathscr{M}}}$ for some $x_{0}\in{\mathscr{M}}_{0}$.
B.1.2. Weight $-1$ and 00
Next, we turn to mixed Hodge structures of weight of $-1$ and 00.
Fix the following data on $M=M_{-1}\oplus{\mathbb{Q}}$: the weight filtration $W_{\bullet}:=(0=W_{-2}M\subseteq W_{-1}M=M_{-1}\subseteq W_{0}M=M)$; the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{M_{-1}}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(M,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
  1. (a)
    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}M_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
  2. (b)
    $\mathrm{Gr}^{W}_{-1}M=M_{-1}$ is polarized by $Q_{-1}$.
See for example [Pea00, below (3.7) to the Remark below Lem. 3.9]. Notice that $\mathrm{Gr}^{W}_{0}M={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$. So $M$ is graded-polarized.
In our case, we need a better understanding of the structure of ${\mathscr{M}}$ than [Pea00]. The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
(B.3) $\mathbf{G}^{{\mathscr{M}}}:=M_{-1}\rtimes\mathrm{Aut}(M_{-1},Q_{-1})=M_{-1}\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
[…]
(B.4) ${\mathscr{M}}=\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}x$
and a $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective
(B.5) ${\mathscr{M}}\longrightarrow{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}),\quad x\mapsto h_{x}.$
As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $M$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$.
B.1.3.
For the quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/M_{-1}$, consider the following surjective $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map
(B.6) $p\colon{\mathscr{M}}\rightarrow{\mathscr{M}}_{0},\quad g\cdot x\mapsto p(g)\cdot\bar{h}_{x}$
with $\bar{h}_{x}=p\circ h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}$. This map sends the mixed Hodge structure $(M,W_{\bullet}M,F^{\bullet}M_{{\mathbb{C}}})$ to the associated pure Hodge structure on $M_{-1}=\mathrm{Gr}_{W}^{-1}M$.
We will study the fibered structure given by $p$ more carefully. We will do this in the finer setting of Mumford–Tate domains.
B.2. Mumford–Tate domains
Definition B.1.
A subset ${\mathscr{D}}$ of the classifying space ${\mathscr{M}}$ is called a (mixed) Mumford–Tate domain if there exists an element $x\in{\mathscr{D}}$ such that ${\mathscr{D}}=\mathbf{G}({\mathbb{R}})^{+}x$, where $\mathbf{G}=\mathrm{MT}(h_{x})$.
The group $\mathbf{G}$ in the definition above is called the generic Mumford–Tate group of ${\mathscr{D}}$ and is denoted by $\mathrm{MT}({\mathscr{D}})$. It is known that $\mathbf{G}<\mathbf{G}^{{\mathscr{M}}}$.
[…]
Recall that ${\mathscr{M}}$ is a semi-algebraic open subset in some algebraic variety ${\mathscr{M}}^{\vee}$ over ${\mathbb{C}}$. Hence, ${\mathscr{D}}$ is a semi-algebraic open subset in some algebraic variety ${\mathscr{D}}^{\vee}$ over ${\mathbb{C}}$. This endows ${\mathscr{D}}$ with a semi-algebraic structure and a complex structure.
Definition B.2.
A subset of ${\mathscr{D}}$ is said to be irreducible algebraic if it is both complex analytic irreducible and semi-algebraic.
In view of [KUY16, Lem. B.1 and its proof], a subset of ${\mathscr{D}}$ is irreducible algebraic if and only if it is a component of $U\cap{\mathscr{D}}$ with $U$ an algebraic subvariety of ${\mathscr{D}}^{\vee}$.
Now, let us take a closer look at the semi-algebraic structure and a complex structure on ${\mathscr{D}}$.
The unipotent radical $V$ of $\mathbf{G}=\mathrm{MT}({\mathscr{D}})$ equals $M_{-1}\cap\mathbf{G}$ by reason of weight. Let $\mathbf{G}_{0}:=\mathbf{G}/V$ be the reductive part. Set ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the map $p$ defined in (B.6). Then ${\mathscr{D}}_{0}$ is a $\mathbf{G}_{0}({\mathbb{R}})^{+}$-orbit and is in fact a (pure) Mumford–Tate domain in the classifying space ${\mathscr{M}}_{0}$, and $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$ for all $x_{0}\in{\mathscr{D}}_{0}$.
By abuse of notation, we also use $p$ to denote the natural projections
(B.7) $p\colon\mathbf{G}\rightarrow\mathbf{G}_{0}=\mathbf{G}/V\quad\text{ and }\quad p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}.$
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Identify $\mathbf{G}_{0}$ with $\{0\}\times\mathbf{G}_{0}$.
Recall that each $x_{0}\in{\mathscr{M}}_{0}$ endows $M_{-1}$ with a Hodge structure of weight $-1$. For $x_{0}\in{\mathscr{D}}_{0}$, $V$ is a sub-Hodge structure because $V$ is a $\mathbf{G}_{0}$-submodule of $M_{-1}$ and that $\mathrm{MT}_{x_{0}}<\mathbf{G}_{0}$.
Finally, consider the constant bundle $V({\mathbb{C}})\times{\mathscr{D}}_{0}\rightarrow{\mathscr{D}}_{0}$. Define the holomorphic subbundle $F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$ to be such that the fiber over each $x_{0}\in{\mathscr{D}}_{0}$ is $F^{0}_{x_{0}}V_{{\mathbb{C}}}$.
Proposition B.3.
The following is true:
  1. (i)
    The semi-algebraic structure on ${\mathscr{D}}\subseteq{\mathrm{Hom}}({\mathbb{S}},\mathbf{G}_{{\mathbb{R}}})$ is given by
    $V({\mathbb{R}})\times{\mathscr{D}}_{0}\xrightarrow{\sim}{\mathscr{D}},\quad(v,x_{0})\mapsto\mathrm{Int}(v)\circ h_{x_{0}}.$
  2. (ii)
    The complex structure on ${\mathscr{D}}$ is given by ${\mathscr{D}}=(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})$.
  3. (iii)
    These two structures are related by the natural bijection
    (B.8) $V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{C}})\times{\mathscr{D}}_{0}\longrightarrow(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0}).$
Proof.
Fix $x\in{\mathscr{D}}$ and let $h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}_{{\mathbb{R}}}$ be the corresponding homomorphism.
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a bijection $\mathbf{G}({\mathbb{R}})^{+}x\cong V({\mathbb{R}})\times{\mathscr{D}}_{0}$, which is semi-algebraic. This establishes (i).
Set ${\mathscr{X}}:=\mathbf{G}({\mathbb{R}})^{+}V({\mathbb{C}})\cdot h_{x}\subseteq{\mathrm{Hom}}({\mathbb{S}}_{{\mathbb{C}}},\mathbf{G}_{{\mathbb{C}}})$, where the action is given by conjugation. The quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}$ induces a natural surjective map ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and by [Pin89, 1.8(a)] each fiber of this map is a $V({\mathbb{C}})$-torsor.
The Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ induces a global section of ${\mathscr{X}}\rightarrow{\mathscr{D}}_{0}$, and hence an isomorphism ${\mathscr{X}}\cong V({\mathbb{C}})\times{\mathscr{D}}_{0}$ over ${\mathscr{D}}_{0}$.
Consider the following surjective equivariant map
(B.9) $\varphi\colon{\mathscr{X}}\rightarrow{\mathscr{D}},\qquad gh_{x}g^{-1}\mapsto g\cdot x.$
By [Pin89, 1.8(b)], for each $x\in{\mathscr{D}}$, the fiber $\varphi^{-1}(x)$ is a principle homogeneous space under $F^{0}_{x_{0}}V_{{\mathbb{C}}}$ with $x_{0}=p(x)$. Hence $\varphi$ gives a bijection $(V({\mathbb{C}})\times{\mathscr{D}}_{0})/F^{0}(V({\mathbb{C}})\times{\mathscr{D}}_{0})\rightarrow{\mathscr{D}}$ over ${\mathscr{D}}_{0}$. Moreover, the complex structure of ${\mathscr{D}}$ is precisely given by this bijection; see [GK24, Proof of Prop. 2.6 in Appendix A] which is a consequence of [Pea00, Thm. 3.13]. This establishes (ii).
[…]
Corollary B.4.
Let $\tilde{Z}_{0}\subseteq{\mathscr{D}}_{0}$ be an irreducible algebraic subset. Then for any $a\in V({\mathbb{R}})$, the subset $\{a\}\times\tilde{Z}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{D}}_{0}\cong{\mathscr{D}}$ is irreducible algebraic.
Proof.
$\{a\}\times\tilde{Z}_{0}$ is clearly semi-algebraic in ${\mathscr{D}}$. In view of (B.8), $\{a\}\times\tilde{Z}_{0}$ is also complex analytic in ${\mathscr{D}}$. Hence, we are done. ∎
Lemma B.5.
Let ${\mathscr{D}}^{\prime}$ be a sub-Mumford–Tate domain of ${\mathscr{D}}$. Then under the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), we have ${\mathscr{D}}^{\prime}=(V^{\prime}({\mathbb{R}})+v_{0})\times p({\mathscr{D}}^{\prime})$ for some $v_{0}\in V({\mathbb{Q}})$.
Proof.
Denote by $\mathbf{G}^{\prime}=\mathrm{MT}({\mathscr{D}}^{\prime})$, $V^{\prime}:=V\cap\mathbf{G}^{\prime}$, and $\mathbf{G}^{\prime}_{0}:=\mathbf{G}^{\prime}/V^{\prime}$. Then, since ${\mathscr{D}}^{\prime}$ is a sub-Mumford–Tate domain of ${\mathscr{D}}$ and by reason of weight, $V^{\prime}$ is the unipotent radical of $\mathbf{G}^{\prime}$.
Set $\mathbf{G}^{\prime\prime}:=V\mathbf{G}^{\prime}$. Then $\mathbf{G}^{\prime\prime}$ is a subgroup of $\mathbf{G}$, with unipotent radical $V$ and reductive part $\mathbf{G}^{\prime}_{0}$. We can construct two Levi decompositions of $\mathbf{G}^{\prime\prime}$ as follows.
First, $\mathbf{G}^{\prime}_{0}=\mathbf{G}^{\prime}/V^{\prime}=\mathbf{G}^{\prime}/(V\cap\mathbf{G}^{\prime})<\mathbf{G}/V=\mathbf{G}_{0}$. Under the Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$ fixed above Proposition B.3, we have $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}$. The composite $\{0\}\times\mathbf{G}^{\prime}_{0}<\mathbf{G}^{\prime\prime}\rightarrow\mathbf{G}^{\prime\prime}/V=\mathbf{G}^{\prime}_{0}$ is the natural isomorphism (the projection to the second factor). Thus, we get a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes\mathbf{G}^{\prime}_{0}$, which is compatible with our fixed Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$.
Next, fix a Levi decomposition $\mathbf{G}^{\prime}=V^{\prime}\rtimes^{\prime}\mathbf{G}^{\prime}_{0}$. It is uniquely determined by a section of $\mathbf{G}^{\prime}\rightarrow\mathbf{G}^{\prime}_{0}$, i.e. an injective morphism $s^{\prime}\colon\mathbf{G}^{\prime}_{0}\rightarrow\mathbf{G}^{\prime}$ such that $s^{\prime}\circ p_{\mathbf{G}^{\prime}}=1_{\mathbf{G}^{\prime}_{0}}$. Now abuse of notation denote by $s^{\prime}$ the composite $\mathbf{G}^{\prime}_{0}\xrightarrow{s^{\prime}}\mathbf{G}^{\prime}<\mathbf{G}^{\prime\prime}$. Then $s^{\prime}$ defines a Levi decomposition $\mathbf{G}^{\prime\prime}=V\rtimes^{\prime}\mathbf{G}_{0}^{\prime}$.
By the general theory of algebraic groups, any two Levi decompositions of $\mathbf{G}^{\prime\prime}$ differ from the conjugation by some $v_{0}\in V({\mathbb{Q}})$. Now, the conclusion follows from Proposition B.3.(i).∎
B.3. Quotient by a normal subgroup and weak Mumford–Tate domains
Let ${\mathscr{D}}$ be a Mumford–Tate domain in ${\mathscr{M}}$ with $\mathrm{MT}({\mathscr{D}})=\mathbf{G}$. Let $N\lhd\mathrm{MT}({\mathscr{D}})$. By [GK24, Prop. 5.1], we have a quotient in the category of complex varieties
$p_{N}\colon{\mathscr{D}}\rightarrow{\mathscr{D}}/N$
with the following properties: (i) ${\mathscr{D}}/N$ is a Mumford–Tate domain in some classifying space of mixed Hodge structures (which must be of weight $-1$ and 00) and $\mathrm{MT}({\mathscr{D}}/N)=\mathbf{G}/N$; (ii) each fiber of $p_{N}$ is an $N({\mathbb{R}})^{+}$-orbit. It is clearly true that $p_{N}$ is semi-algebraic. More precisely, let $M_{-1}/N$ be the maximal quotient of $M_{-1}$ on which $N$ acts trivially, and let $M/N=M_{-1}/N\oplus{\mathbb{Q}}$. For each $x\in{\mathscr{D}}$, the composition
$\bar{h}_{x}:{\mathbb{S}}\overset{h_{x}}{{\longrightarrow}}{\mathrm{GL}}(M_{{\mathbb{R}}}){\longrightarrow}{\mathrm{GL}}((M/N)_{{\mathbb{R}}})$
defines some Hodge structure on $M/N$ with weight $-1$ on $M_{-1}/N$ and weight 00 on the quotient ${\mathbb{Q}}$. The Hodge numbers on $M/N$ defined by $h_{x}$ do not depend on the choice of $x$. Let ${\mathscr{M}}^{\prime}$ denote the corresponding classification space. Then $\bar{h}_{x}$ defined a point $\bar{x}$ in ${\mathscr{M}}^{\prime}$. The images of such $\bar{x}$ form a Mumford–Tate domain denoted by ${\mathscr{D}}/N$. Thus, we have the desired quotient map $p_{N}:{\mathscr{D}}{\longrightarrow}{\mathscr{D}}/N$ which is both complex analytic and real semi-algebraic.
Assume $\Gamma$ is an arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$. Then the quotient $\Gamma\backslash{\mathscr{D}}$ is an orbifold. Denote by $\Gamma_{/N}$ the image of $\Gamma$ under the quotient $\mathbf{G}\rightarrow\mathbf{G}/N$. Then the quotient $p_{N}$ induces
$[p_{N}]\colon\Gamma\backslash{\mathscr{D}}\rightarrow\Gamma_{/N}\backslash({\mathscr{D}}/N).$
We will pay special attention to the fibers of $p_{N}$ (and $[p_{N}]$). More generally, we define:
Definition B.6.
A subset ${\mathscr{D}}_{N}$ of ${\mathscr{D}}$ is called a weak Mumford–Tate domain if there exist $x\in{\mathscr{D}}$ and a normal subgroup $N$ of $\mathrm{MT}(x)$ such that ${\mathscr{D}}_{N}=N({\mathbb{R}})^{+}x$.
In this definition, if $x$ is taken to be a Hodge generic point, i.e. $\mathrm{MT}(x)=\mathbf{G}$, then the weakly Mumford–Tate domain thus obtained is a fiber of $p_{N}$.
Appendix C Intermediate Jacobians and normal functions
[…]
Write ${\mathscr{V}}={\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathscr{O}}_{S}$ for the holomorphic vector bundle obtained from ${\mathbb{V}}_{{\mathbb{Z}}}$.
C.1. Definition and basic property of Intermediate Jacobians
Definition C.1.
The quotient ${\mathscr{V}}/({\mathscr{F}}^{n}+{\mathbb{V}}_{{\mathbb{Z}}})$ is called the (relative) intermediate Jacobian of $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})\rightarrow S$, and is denoted by ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ (or ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ when the Hodge filtration is clear in the context).
Lemma C.2.
$\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ is torus fibration, i.e. a holomorphic family of compact complex torus.
Proof.
For each $s\in S$, we have ${\mathbb{V}}_{{\mathbb{C}},s}={\mathscr{F}}^{n}_{s}\oplus\overline{{\mathscr{F}}^{n}_{s}}$ because ${\mathbb{V}}_{{\mathbb{Z}},s}$ has weight $12n-$ 1. In particular, $\dim{\mathscr{F}}^{n}_{s}=\frac{1}{2}\dim{\mathscr{V}}_{s}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}\cap{\mathscr{F}}^{n}_{s}=\{0\}$. Hence, each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$ is a compact complex torus, and this yields the claim. ∎
Example C.3.
Let us look at the following example from geometry. Let $f\colon X\rightarrow S$ be a smooth projective morphism of relative dimension $d$ over a complex quasi-projective variety such that each fiber is irreducible. For each $n\geq 1$, the relative intermediate Jacobian of the (polarizable) VHS $R^{2n-1}f_{*}{\mathbb{Z}}(n)_{X}$ is called the $n$-th intermediate Jacobian of $X\rightarrow S$ and is denoted by ${\mathscr{J}}^{n}(X/S)$. If $S$ is a point and $X=X$, then we simply write ${\mathscr{J}}^{n}(X)$.
For each $s\in S$ we have $H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{Z}})={\mathbb{V}}_{{\mathbb{Z}},s}$ as ${\mathbb{Z}}$-modules. So the local system $(R^{1}\pi_{*}{\mathbb{Z}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})})^{\vee}$ is ${\mathbb{V}}_{{\mathbb{Z}}}$. In particular, ${\mathscr{V}}_{s}=H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathbb{C}}))$.
The dual of $H^{1,0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\subseteq H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$ gives $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})^{\vee}\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})$. Thus the de Rham–Betti comparison $H^{1}_{\mathrm{dR}}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})\cong H^{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})$ gives rise to a Hodge filtration ${\mathscr{F}}_{{\mathscr{J}}}^{\bullet}$ on ${\mathscr{V}}$ defined by ${\mathscr{F}}_{{\mathscr{J}}}^{1}=0$, ${\mathscr{F}}_{{\mathscr{J}}}^{-1}={\mathscr{V}}$ and
$({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}=\ker({\mathscr{V}}_{s}=H_{1}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s},{\mathbb{C}})\rightarrow\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}))\quad\text{for all }s\in S.$
Lemma C.4.
${\mathscr{F}}_{{\mathscr{J}}}^{0}={\mathscr{F}}^{n}$.
Proof.
We computed in the proof of Lemma C.2 that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{V}}_{s}/({\mathscr{F}}^{n}_{s}+{\mathbb{V}}_{{\mathbb{Z}},s})$. Hence $\mathrm{Lie}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{V}}_{s}/{\mathscr{F}}^{n}_{s}$. So $({\mathscr{F}}_{{\mathscr{J}}}^{0})_{s}={\mathscr{F}}^{n}_{s}$ by defintion. We are done. ∎
By definition, ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}(n))$ and ${\mathbb{V}}_{{\mathbb{Z}}}(n)$ is of weight $-1$. Hence in the rest of the paper, without loss of generality, we assume $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ to have weight $-1$ for the discussion on intermediate Jacobians.
C.2. Betti foliation
Next, we discuss the Betti foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. Write $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ for the natural projection and $d$ for the relative dimension.
The holomorphic vector bundle ${\mathscr{V}}$ is endowed with a connection $\nabla$ whose sheaf of the horizontal sections is ${\mathbb{V}}_{{\mathbb{Z}}}\otimes_{{\mathbb{Z}}}{\mathbb{C}}$.
For each open subset $\Delta\subseteq S$, We have the following exact sequence.
$0\longrightarrow{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}\longrightarrow{\mathscr{V}}_{\Delta}\longrightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)\rightarrow 0$
Locally on $S$, the local system ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}$ is trivial and so is $({\mathscr{V}},\nabla)|_{\Delta}$. Thus ${\mathbb{V}}_{{\mathbb{Z}}}|_{\Delta}\subseteq{\mathscr{V}}_{\Delta}$ becomes ${\mathbb{Z}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$, which extends to ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta$. Notice that $({\mathbb{R}}^{2d}\times\Delta)\cap{\mathscr{F}}_{{\mathscr{J}}}^{0}{\mathscr{V}}_{\Delta}$ is 00 on each fiber by weight reasons. Hence ${\mathbb{R}}^{2d}\times\Delta\subseteq{\mathbb{C}}^{2d}\times\Delta\rightarrow\mathrm{Lie}(\pi^{-1}(\Delta)/\Delta)$ is a real analytic diffeomorphism, and over each $s\in\Delta$ it becomes a group homomorphism. Moreover, the image of $\{r\}\times\Delta$ is easily seen to be complex analytic for any $r\in{\mathbb{R}}^{2d}$.
Taking the quotient of ${\mathbb{V}}_{{\mathbb{Z}}}$ on both sides, we obtain a real analytic diffeomorphism $({\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d})\times\Delta\xrightarrow{\sim}\pi^{-1}(\Delta)$. Let $b_{\Delta}\colon\pi^{-1}(\Delta)\rightarrow{\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d}$ be the composite of its inverse with the projection to the first factor $({\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d})\times\Delta\rightarrow{\mathbb{R}}^{2d}/{\mathbb{Z}}^{2d}$. Then, each fiber of $b_{\Delta}$ is complex analytic by the last sentence of the last paragraph.
The construction above patches to a real analytic homeomorphism
(C.1) ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{\sim}{\mathbb{V}}_{{\mathbb{R}}}/{\mathbb{V}}_{{\mathbb{Z}}}.$
This gives a foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which we call the Betti foliation. More concretely ${\mathscr{F}}_{\mathrm{Betti}}$ is defined as follows: for each $x\in\pi^{-1}(\Delta)$, the local leaf through $x$ is the fiber $b_{\Delta}^{-1}(b_{\Delta}(x))$. Each leaf is holomorphic to the discussion above. In fact, the Betti foliation is the unique foliation on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ which is everywhere transverse to the fibers of $\pi$ and whose set of leaves contains all torsion multisections.
For each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, the Betti foliation induces a decomposition $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$, and $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ induces a natural isomorphism $T_{x}{\mathscr{F}}_{\mathrm{Betti}}\cong T_{\pi(x)}S$. The translation on the torus ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{\pi(x)}$ yields a canonical isomorphism $T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})$. Hence, we have a linear map
(C.2) $q_{x}\colon T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})=T_{x}{\mathscr{F}}_{\mathrm{Betti}}\oplus T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})\rightarrow T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)})=T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)}),$
whose kernel is $T_{x}{\mathscr{F}}_{\mathrm{Betti}}$.
We close this subsection with the following discussion. Take a holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ of the intermediate Jacobian. Then we have a linear map, at each $s\in S$,
(C.3) $\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{q_{\nu(s)}}T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s}).$
Notice that, since $\pi$ induces canonically $T_{\nu(s)}{\mathscr{F}}_{\mathrm{Betti}}=T_{s}S$, the map $\mathrm{d}\nu$ is exactly $(1,\nu_{\mathrm{Betti},s})\colon T_{s}S\rightarrow T_{s}S\oplus T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})=T_{s}S\oplus T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})=T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$.
C.3. Sections of intermediate Jacobians
In this subsection, we explain how any holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ defines a family of mixed Hodge structures on $S$ which varies holomorphically (Definition A.4 without the Griffiths’ transversality) of weight $-1$ and 00.
First we define a local system ${\mathbb{E}}_{\nu}$ on $S$ associated with $\nu$. Write ${\mathscr{J}}$ for the sheaf of holomorphic sections of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Then we have the following exact sequence
$0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathscr{V}}/{\mathscr{F}}^{0}{\mathscr{V}}\rightarrow{\mathscr{J}}\rightarrow 0$
Taking cohomology yields the boundary map.
$c\colon H^{0}(S,{\mathscr{J}})\rightarrow H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}}).$
On the other hand, it is known that $H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}})$ can be canonically identified with $\mathrm{Ext}_{\mathrm{loc.sys}}({\mathbb{Z}}_{S},{\mathbb{V}}_{{\mathbb{Z}}})$, the isomorphism classes of the extensions of local systems (${\mathbb{Z}}_{S}$ by ${\mathbb{V}}_{{\mathbb{Z}}}$) on $S$, so $c(\nu)\in H^{1}(S,{\mathbb{V}}_{{\mathbb{Z}}})$ defines a local system ${\mathbb{E}}_{\nu}$ on $S$ fitting into the short exact sequence $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}_{S}\rightarrow 0$. Notice that this defines a weight filtration $W_{\bullet}$ on ${\mathbb{E}}_{\nu}$ of weight $-1$ and 00, by letting $W_{-2}{\mathbb{E}}_{\nu}=0$ and $W_{-1}{\mathbb{E}}_{\nu}={\mathbb{V}}_{{\mathbb{Z}}}$ and $W_{0}{\mathbb{E}}_{\nu}={\mathbb{E}}_{\nu}$. Notice that in the category of local systems, this exact sequence is split after $\otimes{\mathbb{Q}}$ if $S$ is simply connected.
Carlson [Car85] proved that ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$, the set of congruence classes of extensions of ${\mathbb{Z}}(0)$ by ${\mathbb{V}}_{{\mathbb{Z}},s}$ in the category of mixed Hodge structures; see for example [BZ14, Thm. 8.4.2]. In our context, this says the following. For each $s\in S({\mathbb{C}})$, $\nu(s)$ defines a mixed Hodge structure on ${\mathbb{E}}_{\nu,s}$, and hence a Hodge filtration ${\mathscr{F}}_{{\mathbb{E}}}^{\bullet}$ on the fiber ${\mathbb{E}}_{\nu,s}\otimes{\mathbb{C}}$. Since $\nu$ is holomorphic, the fiberwise Hodge filtrations give rise to a Hodge filtration on ${\mathbb{E}}_{\nu}$. So $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ is a family of mixed Hodge structures on $S$ which varies holomorphically and has weight $-1$ and 00. We are done.
Moreover, if ${\mathbb{V}}_{{\mathbb{Z}}}$ is polarized, then ${\mathbb{E}}_{\nu}$ is graded-polarized with this polarization on ${\mathbb{V}}_{{\mathbb{Z}}}$ and the canonical polarization on ${\mathbb{Z}}(0)$ given by ${\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$, $a\otimes b\mapsto ab$.
C.4. Normal functions
[…]
Definition C.5.
A holomorphic section $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is called an admissible normal function if $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ is an admissible VMHS.
C.5. Normal functions arising from families of algebraic cycles
Let $X$ be a smooth irreducible projective variety over ${\mathbb{C}}$ with $\dim X=d$. For each $n$, the Chow group of $n$-cocycles $\mathrm{CH}^{n}(X)$ is the group of algebraic cycles of codimension $n$ on $X$ modulo rational equivalence. Denote by $\mathrm{CH}^{n}(X)_{\hom}$ the kernel of the cycle map $\mathrm{CH}^{n}(X)\rightarrow H^{2n}(X,{\mathbb{Z}})$. An $n$-cocycle $Z$ is said to be homologically trivial if its Chow class lies in $\mathrm{CH}^{n}(X)_{\hom}$. By abuse of notation, we will denote the Chow class of a (co)cycle $Z$ also by $Z$.
Let ${\mathscr{J}}^{n}(X)$ be the intermediate Jacobian corresponding to $H^{2n-1}(X,{\mathbb{Z}})$ from Example C.3.
The Abel–Jacobi map
(C.4) ${\mathrm{AJ}}\colon\mathrm{CH}^{n}(X)_{\hom}\longrightarrow{\mathscr{J}}^{n}(X)$
[…]
Now, we turn to the family version and define the corresponding normal function. Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $Z$ be a family of homologically trivial $n$-cocycle in $X/S$, i.e. a formal sum of integral subschemes of $X$ which are flat and dominant over $S$ such that each fiber $Z_{s}$ is a homologically trivial $n$-cocycle of $X_{s}$.
Theorem C.6 ([EZ86]).
The holomorphic section
$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{n}(X/S),\qquad s\mapsto{\mathrm{AJ}}(Z_{s})$
It is an admissible normal function.
The following proposition is a simple application of the Abel–Jacobi map and the Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on intermediate Jacobians introduced in §$\S$§C.2.
Proposition C.7.
Set $S^{\circ}:=S\setminus S_{{\mathscr{F}}}(1)$.
Assume $X/S$ and $Z$ are defined over $\overline{\mathbb{Q}}$. Then $[Z_{s}]\in\mathrm{CH}^{n}(X_{s})$ is non-torsion for any $s\in S^{\circ}({\mathbb{C}})\setminus S^{\circ}(\overline{\mathbb{Q}})$.
Proof.
Take $s\in S({\mathbb{C}})$, and set $\bar{s}$ to be the $\overline{\mathbb{Q}}$-Zariski closure of $s$ in $S$.
Assume $[Z_{s}]$ is a torsion point of $\mathrm{CH}^{n}(X_{s})$. Then $[Z_{t}]$ is a torsion point of $\mathrm{CH}^{n}(X_{t})$ for each $t\in\bar{s}(\overline{\mathbb{Q}})$, and $t\mapsto{\mathrm{AJ}}(Z_{t})$ is a torsion section of ${\mathscr{J}}^{n}(X/S)\times_{S}\bar{s}\rightarrow\bar{s}$. So $\nu_{Z}(\bar{s})\subseteq{\mathscr{F}}_{\mathrm{Betti}}$. If $s\not\in S(\overline{\mathbb{Q}})$, then $\dim\bar{s}\geq 1$, and hence $\bar{s}\subseteq S_{{\mathscr{F}}}(1)$ by definition. Now we are done. ∎
C.6. Normal functions associated with Gross–Schoen cycles and Ceresa cycles
Let $C$ be an irreducible smooth projective curve defined over a field $k$ of genus $g\geq 3$. Let $\xi\in C(k)$. We have:
    For each subset $T\subseteq\{1,2,3\}$, the modified diagonal $\Delta_{T}(C)=\{(x_{1},x_{2},x_{3}):x_{i}=\xi\text{ for }i\not\in T,~{}x_{j}=x_{j^{\prime}}\text{ for all }j,j^{\prime}\in T\}\in\mathrm{CH}^{2}(C^{3})$.
    The (classical) Abel–Jacobi map $i_{\xi}\colon C\rightarrow\mathrm{Jac}(C)$ sending $x\mapsto[x-\xi]$.
Both $\Delta_{T}(C)$ and $i_{\xi}$ can be extended to any $\xi=\sum_{i}n_{i}e_{i}\in\mathrm{Div}^{1}(C)(k)=\mathrm{Pic}^{1}(C)(k)$. This is classical and direct for $i_{\xi}$, and for $\Delta_{T}(C)$ one can define $\Delta_{123}(C)=\{(c,c,c):c\in C\}$, $\Delta_{12}(C)=\sum n_{i}\{(c,c,e_{i}):c\in C\}$, $\Delta_{1}(C)=\sum_{i,j}n_{i}n_{j}\{(c,e_{i},e_{j}):c\in C\}$, etc.
For any $\xi\in\mathrm{Pic}^{1}(C)$, define the Gross–Schoen cycle based at $\xi$ (resp. the Ceresa cycle based at $\xi$) to be:
  • (Gross–Schoen cycle based at $\xi$) $\Delta_{{\mathrm{GS}},\xi}(C):=\Delta_{123}(C)-\Delta_{12}(C)-\Delta_{13}(C)-\Delta_{23}(C)+\Delta_{1}(C)+\Delta_{2}(C)+\Delta_{3}(C)\in\mathrm{CH}^{2}(C^{3})$.
  • (Ceresa cycle based at $\xi$) ${\mathrm{Ce}}_{\xi}(C):=i_{\xi}(C)-[-1]_{*}i_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$.
${\mathrm{Ce}}_{\xi}(C)$ is clearly homologically trivial because $[-1]$ acts trivially on even-degree cohomology groups. It is not hard to check that $\Delta_{{\mathrm{GS}},\xi}(C)$ is also homologically trivial; this follows for example from [GS95, Prop. 3.1] and because the map $\mathrm{Pic}^{1}(C)\rightarrow\mathrm{CH}^{2}(C^{3})$, $\xi\mapsto\Delta_{{\mathrm{GS}},\xi}(C)$ is a group homomorphism.
Let $\omega_{C}$ be the canonical divisor on $C$.
Definition C.8.
Let $\xi\in\mathrm{Pic}^{1}(C)(k)$ such that $(2g-2)\xi=\omega_{C}$. Define the (canonical) Gross–Schoen cycle and the (canonical) Ceresa cycle of $C$ to be:
  • (Gross–Schoen cycle) $\Delta_{{\mathrm{GS}}}(C):=\Delta_{{\mathrm{GS}},\xi}(C)\in\mathrm{CH}^{2}(C^{3})_{\hom}$.
  • (Ceresa cycle) ${\mathrm{Ce}}(C):={\mathrm{Ce}}_{\xi}(C)\in\mathrm{CH}^{g-1}(\mathrm{Jac}(C))_{\hom}$.
There are finitely many $\xi$’s with $(2g-2)\xi=\omega_{C}$ and each two differ from a $(2g-2)$-torsion. Hence, the Gross–Schoen cycle and the Ceresa cycle of $C$ are well-defined at $(2g-2)$-torsions.
We will associate normal functions with the Schoen cycles and Ceresa cycles. First we have ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}^{2}(C^{3})={\mathscr{J}}(H^{3}(C^{3},{\mathbb{Z}})(2))$ and ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}^{g-1}(\mathrm{Jac}(C))={\mathscr{J}}(H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1))$. Using the Poincaré duality on $\mathrm{Jac}(C)$ and on $C$, we get $H^{2g-3}(\mathrm{Jac}(C),{\mathbb{Z}})(g-1)=H_{3}(\mathrm{Jac}(C),{\mathbb{Z}})(-1)=\bigwedge^{3}H_{1}(C,{\mathbb{Z}})(-1)=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)$. So ${\mathrm{AJ}}({\mathrm{Ce}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
We have $\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2)\subseteq H^{3}(C^{3},{\mathbb{Z}})(2)$ in the following way. First, the Künneth formula gives a decomposition $H^{3}(C^{3},{\mathbb{Z}})=H^{1}(C,{\mathbb{Z}})^{\otimes 3}\bigoplus H^{1}(C,{\mathbb{Z}})(-1)^{\oplus 6}$. Next, $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$, as a subspace of $H^{3}(C^{3},{\mathbb{Z}})$, has a basis consisting of $\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3}$, with $\alpha_{j}$ the pullback of an element in $H^{1}(C,{\mathbb{Z}})$ under the $j$-th projection $C^{3}\rightarrow C$. The symmetric group $S_{3}$ acts naturally on $C^{3}$, and this induces an action of $S_{3}$ on $H^{1}(C,{\mathbb{Z}})^{\otimes 3}$ with $\sigma(\alpha_{1}\wedge\alpha_{2}\wedge\alpha_{3})=\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$ for each $\sigma\in S_{3}$. Then $(H^{1}(C,{\mathbb{Z}})^{\otimes 3})^{S_{3}}=\bigwedge^{3}H^{1}(C,{\mathbb{Z}})$ for this action, with each member having a basis consisting of $\sum_{\sigma\in S_{3}}\mathrm{sgn}(\sigma)\alpha_{\sigma(1)}\wedge\alpha_{\sigma(2)}\wedge\alpha_{\sigma(3)}$.
It is easy to check that the pushforward of $\Delta_{{\mathrm{GS}}}(C)$ to any two factors of $C^{3}$ is trivial. Hence ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}(H^{1}(C,{\mathbb{Z}})^{\otimes 3}(2))$. Moreover, the modified diagonal is easily seen to be invariant under the action of $S_{3}$ on $C^{3}$. Thus ${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))\in{\mathscr{J}}(\bigwedge^{3}H^{1}(C,{\mathbb{Z}})(2))$.
Lemma C.9.
${\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}(C))=3{\mathrm{AJ}}({\mathrm{Ce}}(C))$.
Proof.
Consider $\iota_{3}\colon C^{3}\rightarrow\mathrm{Jac}(C)$, $(c_{1},c_{2},c_{3})\mapsto i_{\xi}(c_{1})+i_{\xi}(c_{2})+i_{\xi}(c_{3})$. The difference $(\iota_{3})_{*}\Delta_{{\mathrm{GS}}}(C)-3{\mathrm{Ce}}(C)$ was computed by the second-named author [Zha10, Thm. 1.5.5]. More precisely, the Fourier–Mukai transformation yields a spectrum decomposition $C=\sum_{j=0}^{g-1}C_{j}$ in $\mathrm{CH}^{g-1}(\mathrm{Jac}(C))$, with $[n]_{*}C_{j}=n^{2+j}C_{j}$ for all $n\in{\mathbb{Z}}$ and $j\in\{0,\ldots,g-1\}$, and [Zha10, Thm. 1.5.5 and its proof] implies that
$(\iota_{3})_{*}\Delta_{{\mathrm{GS}}}(C)-3{\mathrm{Ce}}(C)=\sum\nolimits_{j\geq 2}a_{j}C_{j}$
For appropriate numbers $a_{j}$. The multiplication $[n]\colon\mathrm{Jac}(C)\rightarrow\mathrm{Jac}(C)$ induces $\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})\rightarrow\bigwedge\nolimits^{3}H^{1}(C,{\mathbb{Z}})$, $x\mapsto n^{3}x$. Hence $n^{2+j}{\mathrm{AJ}}(C_{j})=n^{3}{\mathrm{AJ}}(C_{j})$ for all $n\in{\mathbb{Z}}$, and so ${\mathrm{AJ}}(C_{j})=0$ for all $j\geq 2$.
Hence the conclusion follows because the induced map of $\iota_{3}$ is the natural projection.∎
Now, we turn to defining and studying the normal functions. Let ${\mathscr{M}}_{g}$ be the moduli space of curves of genus $g$, and let $f\colon{\mathscr{C}}_{g}\rightarrow{\mathscr{M}}_{g}$ be the universal curve. We have a VHS $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}\rightarrow{\mathscr{M}}_{g}$. The fiberwise polarization $q_{H}\colon H^{1}(C,{\mathbb{Z}})\otimes H^{1}(C,{\mathbb{Z}})\xrightarrow{\cup}H^{2}(C,{\mathbb{Z}})\cong{\mathbb{Z}}(-1)$, with $\cup$ the cup product (the dual of the intersection pairing) on $H^{1}(C,{\mathbb{Z}})$, induces a polarization on $\bigwedge^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}$.
Now we have two normal functions
(C.5) $\nu_{{\mathrm{GS}}}\colon\mathbb{M}_{g}\longrightarrow{\mathscr{J}}(\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}),$ $\qquad s\mapsto{\mathrm{AJ}}(\Delta_{{\mathrm{GS}}}({\mathscr{C}}_{s})),$
$\nu_{{\mathrm{Ce}}}\colon\mathbb{M}_{g}\longrightarrow{\mathscr{J}}(\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}),$ $\qquad s\mapsto{\mathrm{AJ}}({\mathrm{Ce}}({\mathscr{C}}_{s})).$
We can do better.
Proposition C.10.
Let ${\mathbb{V}}_{{\mathbb{Z}}}$ be the kernel of the morphism of VHS (called the contractor)
(C.6) $c\colon\bigwedge\nolimits^{3}R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}(2)\longrightarrow R^{1}f_{*}{\mathbb{Z}}_{{\mathscr{C}}_{g}}(1)$
fiberwise defined by $x\wedge y\wedge z\mapsto q_{H}(y,z)x+q_{H}(z,x)y+q_{H}(x,y)z$. Then
  1. (i)
    both $\nu_{{\mathrm{GS}}}$ and $\nu_{{\mathrm{Ce}}}$ have images in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$;
  2. (ii)
    neither $\nu_{{\mathrm{GS}}}$ nor $\nu_{{\mathrm{Ce}}}$ is a torsion section;
  3. (iii)
    ${\mathbb{V}}_{{\mathbb{Z}}}$ is an irreducible VHS on ${\mathscr{M}}_{g}$, i.e. the only sub-VHSs of ${\mathbb{V}}_{{\mathbb{Z}}}$ are 00 and itself.
Proof.
For each $s\in S({\mathbb{C}})$, $\nu_{{\mathrm{GS}}}(s)\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ by [Zha10, Lem. 5.1.5], and hence $\nu_{{\mathrm{Ce}}}(s)\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ by Lemma C.9. So (i) holds.
For (ii), it suffices to find a curve $C$ of genus $g$ such that ${\mathrm{AJ}}({\mathrm{Ce}}(C))$ is not torsion. There are many examples of such curves in existing literature. Alternatively, the result for $\nu_{{\mathrm{Ce}}}$ can be already deduced from Ceresa’s original argument in [Cer83].
For (iii), it suffices to prove that $V:={\mathbb{V}}_{{\mathbb{Z}},s}\otimes{\mathbb{Q}}$ is a simple $\mathrm{Sp}_{2g}$-module for one (and hence all) $s\in{\mathscr{M}}_{g}({\mathbb{C}})$, or equivalently $V$ is an irreducible representation of $\mathrm{Sp}_{2g}$. This is a standard result of the representation theory for $\mathfrak{sp}_{2g}$, see for example [FH13, Thm. 17.5]. ∎
Appendix D Metrized Poincaré bundle and local height pairing
Let $S$ be a smooth irreducible quasi-projective variety.
D.1. Height pairing at Archimedean places
Let $f\colon X\rightarrow S$ be a smooth projective morphism of algebraic varieties with irreducible fibers of dimension $d$. Let $p,q$ be non-negative integers such that $p+q=d+1$.
Let $Z$, resp. $W$, be a family of homologically trivial $p$-cocycle in $X/S$, resp. a family of homologically trivial $q$-cocycle in $X/S$. Assume $Z$ and $W$ have disjoint supports over the generic fiber, then up to replacing $S$ by a Zariski open dense subset, we may assume that $Z_{s}$ and $W_{s}$ have disjoint supports for all $s\in S({\mathbb{C}})$.
In Arakelov’s theory, the local height pairing at archimedean places in this context is defined by
(D.1) $\langle Z_{s},W_{s}\rangle_{\infty}=\int_{Z_{s}}\eta_{W_{s}}.$
where $\eta_{W_{s}}$ is a Green’s current for $W_{s}$. This pairing is known to be symmetric.
The two families of homologically trivial cocycles give rise to two admissible normal functions.
$\nu_{Z}\colon S\rightarrow{\mathscr{J}}^{p}(X/S)\quad\text{ and }\quad\nu_{W}\colon S\rightarrow{\mathscr{J}}^{q}(X/S).$
A particularly important case is when $p=q=\frac{d+1}{2}$ and that $Z$ and $W$ are rationally equivalent. In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the height of $Z_{s}$. More precisely, let $Z$ be a family of homologically trivial $p$-cocycle in $X/S$ with $p=\frac{d+1}{2}$, and let $\eta$ be the generic point of $S$. By Moving Lemma, there exists a $p$-cocycle $W_{\eta}$ of $X_{\eta}$ whose class in $\mathrm{CH}^{p}(X_{\eta})$ is the same as the class of $Z_{\eta}$ such that $W_{\eta}$ and $Z_{\eta}$ have disjoint supports. Then $W_{\eta}$ extends to a family of homologically trivial $p$-cocycle $W$ in $X/S$, and $W_{s}$ and $Z_{s}$ have disjoint supports for $s$ in a Zariski open dense subset of $S$. Then we set $\langle Z_{s},Z_{s}\rangle_{\infty}=\langle Z_{s},W_{s}\rangle_{\infty}$. Notice that in this case $\nu_{Z}=\nu_{W}$.
Hain [Hai90, §$\S$3.3] related the local height pairing at archimedean places (D.1) to the metrized Poincaré bundle. We will review this in the next subsection.
D.2. Metrized Poincaré bundle
Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a polarized VHS on $S$ of weight $-1$. The intermediate Jacobian ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is a torus fibration by Lemma C.2. The dual torus fibration $\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}))$ can be described as follows. Set ${\mathbb{V}}_{{\mathbb{Z}}}^{\vee}:={\mathrm{Hom}}_{\mathrm{VHS}}({\mathbb{V}}_{{\mathbb{Z}}},{\mathbb{Z}}(1)_{S})$ with the natural Hodge filtration. Then ${\mathbb{V}}_{{\mathbb{Z}}}^{\vee}$ is a VHS on $S$ of weight $-1$, and there is a canonical isomorphism ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})=\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}))$.
The general theory of biextension says that there exists a unique line bundle ${\mathscr{P}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})$ satisfying the following properties:
  1. (i)
    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\cong{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
  2. (ii)
    over each $s\in S({\mathbb{C}})$, ${\mathscr{P}}|_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times\{\lambda\}}$ represents $\lambda\in\mathrm{Pic}^{0}({\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s})={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}_{s}$,
  3. (iii)
    $\epsilon^{*}{\mathscr{P}}\cong{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\cong({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
Definition D.1.
The metrized line bundle $\overline{{\mathscr{P}}}:=({\mathscr{P}},\|\cdot\|_{\mathrm{can}})$ is called the metrized Poincaré bundle.
[…]
Going from ${\mathbb{Z}}$ to ${\mathbb{R}}$-coefficients, one has the set ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})$ of mixed ${\mathbb{R}}$-Hodge structures of weight $0,-1,-2$ whose weight graded pieces are ${\mathbb{R}}(0)$, ${\mathbb{V}}_{{\mathbb{R}},s}$ and ${\mathbb{R}}(1)$. One can check that ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})$ is canonically isomorphic to ${\mathbb{R}}$. Then on ${\mathscr{P}}^{*}_{s}={\mathscr{B}}({\mathbb{V}}_{{\mathbb{Z}},s})$,
$\|p\|_{\mathrm{can}}=e^{\mathrm{f}_{s}(p)}$
where $\mathrm{f}_{s}$ is the forgetful map ${\mathscr{B}}({\mathbb{V}}_{{\mathbb{Z}},s})\rightarrow{\mathscr{B}}({\mathbb{V}}_{{\mathbb{R}},s})={\mathbb{R}}$.
Now, let us go back to the setting of §$\S$§D.1. Recall the section $(\nu_{Z},\nu_{W})\colon S\rightarrow{\mathscr{J}}^{p}(X/S)\times_{S}{\mathscr{J}}^{q}(X/S)$ obtained from the two families of homologically trivial cocycles $Z$ and $W$.
[…]
Hain [Hai90, Prop. 3.3.2] constructed a section $\beta_{Z,W}$ of the line bundle $(\nu_{Z},\nu_{W})^{*}{\mathscr{P}}\rightarrow S$ in view of the Hodge-theoretic construction of ${\mathscr{P}}$ explained below Definition D.1, and proved [Hai90, Prop. 3.3.12] $\log\|\beta_{Z,W}(s)\|=-\int_{Z_{s}}\eta_{W_{s}}$ for all $s\in S({\mathbb{C}})$. To summarize, we have
Proposition D.2 (Hain).
We have $-\log\|\beta_{Z,W}(s)\|=\langle Z_{s},W_{s}\rangle_{\infty}$ for all $s\in S({\mathbb{C}})$.
D.3. Metrized tautological bundle
As explained at the end of §$\S$§D.1, we are often more interested in the case where ${\mathscr{J}}^{p}(X/S)$ is self-dual and $\nu_{Z}=\nu_{W}$. In this case, $(\nu_{Z},\nu_{W})^{*}\overline{{\mathscr{P}}}=\nu_{Z}^{*}\Delta^{*}\overline{{\mathscr{P}}}$ for the diagonal $\Delta\colon{\mathscr{J}}^{p}(X/S)\rightarrow{\mathscr{J}}^{p}(X/S)\times_{S}{\mathscr{J}}^{p}(X/S)$. We discuss this case in this and the next subsections in a more general setting.
Let $({\mathbb{V}}_{{\mathbb{Z}}},{\mathscr{F}}^{\bullet})$ be a VHS on $S$ of weight $-1$, with a polarization ${\mathscr{Q}}\colon{\mathbb{V}}_{{\mathbb{Z}}}\otimes{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{Z}}(1)_{S}$. Then ${\mathscr{Q}}$ induces a morphism of VHS ${\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}^{\vee}$, and hence a morphism between the intermediate Jacobians $i_{{\mathscr{Q}}}\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\longrightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}$. We thus have a morphism
$\Delta_{{\mathscr{Q}}}=(1,i_{{\mathscr{Q}}})\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}.$
This is the case, for example, for Poincaré duality, where the polarization is given by the cup product.
Definition D.3.
The metrized line bundle $\overline{{\mathscr{P}}^{\Delta_{\mathscr{Q}}}}:=\Delta_{{\mathscr{Q}}}^{*}\overline{{\mathscr{P}}}$ is called the metrized tautological bundle on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$. When the polarization ${\mathscr{Q}}$ is clear in context, we simply denote it by $\overline{{\mathscr{P}}^{\Delta}}$.
By [HR04, Prop. 7.1 and 7.3], the curvature form $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is a closed $2$-form, uniquely determined by the following properties.
Proposition D.4.
[…]
  1. (i)
    Restricted to each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}$, $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is the unique translation invariant $2$-form $\omega_{s}$ given by $2{\mathscr{Q}}_{s}$,
  2. (ii)
    $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ vanishes along each leaf of the Betti foliation.
Here is a more explicit formula for $c_{1}(\overline{{\mathscr{P}}^{\Delta}})$. Write $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ for the natural projection. Recall the linear map (C.2)
$q_{x}\colon T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},\pi(x)}).$
with kernel $T_{x}{\mathscr{F}}_{\mathrm{Betti}}$. Now for any $v_{1},v_{2}\in T_{x}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$, we have
(D.2) $c_{1}(\overline{{\mathscr{P}}^{\Delta}})(v_{1},v_{2})=2{\mathscr{Q}}_{\pi(x)}(q_{x}(v_{1}),q_{x}(v_{2})).$
D.4. Pullback by admissible normal functions
Retain the setup in §$\S$§D.3. Now we turn to admissible normal functions $\nu\colon S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$.
Hain proved [Hai13, Thm. 13.1] that $\nu^{*}c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is a semi-positive $(1,1)$-form.
Definition D.5.
The semi-positive $(1,1)$-form $\nu^{*}c_{1}(\overline{{\mathscr{P}}^{\Delta}})$ is called the Betti form associated with $\nu$, which we denote by $\beta_{\nu}$.
The Betti foliation ${\mathscr{F}}_{\mathrm{Betti}}$ on ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ induces a linear map (C.3) at each $s\in S({\mathbb{C}})$
$\nu_{\mathrm{Betti},s}\colon T_{s}S\xrightarrow{\mathrm{d}\nu}T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\xrightarrow{q_{\nu(s)}}T_{0}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s}).$
By (D.2), $\beta_{\nu}(u,\bar{u})=2{\mathscr{Q}}_{s}(\nu_{\mathrm{Betti},s}(u),\overline{\nu_{\mathrm{Betti},s}(u)})$ for any $u\in T_{s}S$. By Griffiths’ transversality, $\nu_{\mathrm{Betti},s}(u)\in{\mathbb{V}}_{s}^{-1,0}$. So
(D.3) $\beta_{\nu}(u,\bar{u})\geq 0\text{ for all }u\in T_{s}S,\text{ with equality if and only if }\nu_{\mathrm{Betti},s}(u)=0.$
Notice that this also explains the semi-positivity of $\beta_{\nu}$. Now we use (D.3) to prove the following proposition.
Proposition D.6.
For any $s\in S({\mathbb{C}})$, the following are equivalent:
  1. (i)
    $(\beta_{\nu}^{\wedge\dim S})_{s}\not=0$;
  2. (ii)
    $\dim\nu_{\mathrm{Betti},s}(T_{s}S)=\dim S$.
This proposition has the following immediate corollary.
Corollary D.7.
$\beta_{\nu}^{\wedge\dim S}\not\equiv 0$ if and only if $\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S)=\dim S$.
We will call $\max_{s\in S({\mathbb{C}})}\dim\nu_{\mathrm{Betti},s}(T_{s}S)$ the Betti rank of $\nu$ and denote it by $r(\nu)$.
Proof of Proposition D.6.
Assume (i) is false, i.e. $(\beta_{\nu}^{\wedge\dim S})_{s}=0$. By (D.3) there exists $0\not=u\in T_{s}S$ with $\nu_{\mathrm{Betti},s}(u)=0$. Thus $\ker\nu_{\mathrm{Betti},s}\not=0$, and therefore $\dim\nu_{\mathrm{Betti},s}(T_{s}S)<\dim S$. So (ii) is also false.
Assume (ii) is false. Then there exists $0\not=u\in\ker\nu_{\mathrm{Betti},s}$. By (D.3), $\beta_{\nu}(u,\bar{u})=0$. Thus $u$ is an eigenvector of the Hermitian matrix defining $\beta_{\nu}$ with eigenvalue 00. Hence the determinant of this matrix is 00, so $\beta_{\nu}^{\wedge\dim S}=0$ at $s$. So (i) is also false. ∎